### v1

# Data-Driven Design Rules for TADF Emitters from a High-Throughput Screening of $747$ Molecules

Abstract
[…]
I Introduction
Organic Light-Emitting Diodes (OLEDs) based on Thermally Activated Delayed Fluorescence (TADF) have defined the third generation of electroluminescent materials, offering a viable path to achieving near-unity internal quantum efficiency (IQE) without resorting to the expensive and scarce noble metals found in phosphorescent emitters [1, 2, 3]. The TADF mechanism hinges on the efficient harvesting of non-emissive triplet excitons ($T_{1}$) by up-converting them to the emissive singlet state ($S_{1}$) through Reverse Intersystem Crossing (RISC). This process requires satisfying both a thermodynamic condition—a minimized singlet-triplet energy gap ($\Delta E_{\text{ST}}\qty$$\lesssim\qty{0.2}{}$)—and a kinetic condition: a sufficiently high rate of RISC ($k_{\text{RISC}}$) [4, 5].
The rational design of high-performance TADF emitters is a formidable challenge, dominated by a fundamental trade-off. In conventional Donor-Acceptor (D-A) architectures, achieving a small $\Delta E_{\text{ST}}$necessitates a significant spatial separation of the frontier molecular orbitals to minimize the electron exchange energy [6]. However, this separation inherently reduces the orbital overlap, which not only suppresses the oscillator strength ($f$) and thus the radiative decay rate, but can also hinder the spin-orbit coupling (SOC) necessary for efficient RISC [7, 8]. Successfully navigating this delicate balance requires exploring a vast and complex chemical space [9].
Computational chemistry is an indispensable tool for navigating this landscape, yet it presents its own dilemma. High-fidelity methods are computationally prohibitive for screening the thousands of potential candidates required for data-driven discovery [10, 11, 12]. This scalability bottleneck has historically limited computational studies to small, homologous series of molecules, hindering the extraction of universal design principles.
To bridge this gap, we employ a high-throughput virtual screening (HTVS) protocol based on accelerated semi-empirical methods, which we rigorously validated on a diverse dataset of $747$ experimentally known TADF emitters in our preceding companion paper (hereafter, Article 1) [13]. That study established our framework as a powerful and validated tool for large-scale molecular ranking and trend analysis.
In this article, we leverage our validated HTVS protocol to move beyond benchmarking and perform a large-scale, data-driven analysis aimed at elucidating the fundamental structure-property relationships that govern TADF performance. By analyzing the results from the entire 747-molecule dataset, we aim to distill the complex interplay of molecular structure, conformation, and electronic properties into a set of clear, quantitative, and actionable design guidelines. This study, therefore, seeks to: (i) establish a quantitative performance hierarchy among diverse molecular architectures (D-A, D-A-D, and MR systems); (ii) derive statistically robust design rules linking key structural parameters to both thermodynamic ($\Delta E_{\text{ST}}$) and kinetic (SOC, $k_{\text{RISC}}$) factors; (iii) identify and characterize distinct molecular families through data-driven clustering analysis; and (iv) propose a curated list of high-priority candidate molecules predicted to exhibit exceptional TADF properties. By translating a massive computational dataset into actionable design intelligence, this work provides a unified framework to guide and accelerate the discovery of next-generation TADF emitters.
This paper is structured as follows: Section II details our methodological framework. The main results are presented in the subsequent sections, including an analysis of the architectural landscape, a classification of molecular families, and a deep dive into the MR-TADF design paradigm. Finally, Section IV summarizes our conclusions and proposes future directions for TADF emitter design.
II Methodological framework for large-scale analysis
All structure-property relationships presented in this work were derived from a large-scale dataset of $747$ experimentally known TADF emitters, processed using a consistent and validated high-throughput virtual screening (HTVS) protocol. The full details, validation, and limitations of this computational framework are established in our preceding work (Article 1) [13]. Here, we provide a concise summary of the essential methodological components and their justifications.
II.1 Dataset and molecular structures
The dataset comprises a diverse set of $747$ TADF molecules curated from the literature [14]. To facilitate systematic analysis, molecules were classified into distinct architectural families (e.g., D-A, D-A-D, Multi-Resonance) using automated SMARTS pattern matching, followed by manual verification.
Initial 3D structures were generated from SMILES strings, and a rigorous conformational search was performed for each molecule using the CREST program coupled with the GFN2-xTB semi-empirical Hamiltonian [15, 16]. The lowest-energy conformer was then fully optimized at the GFN2-xTB level to obtain the final ground-state ($S_{0}$) geometry. This tight-binding method is specifically parameterized for accurate molecular structures and noncovalent interactions, making it well-suited for the complex and often flexible systems studied here [17].
II.2 Methodological approximations and justifications
II.2.1 Hybrid protocol and the vertical approximation compromise
The HTVS protocol relies on a hybrid computational strategy: ground-state ($S_{0}$) geometries are optimized using GFN2-xTB [16, 18], while excited-state properties ($S_{1}$, $T_{1}$) are calculated using the simplified methods, sTDA-xTB/sTD-DFT-xTB [19, 20]. This methodology implicitly employs the vertical approximation.
[…]
II.2.2 Limitations of the solvation model
The study of excited-state properties was performed using the ground-state-based ALPB implicit solvation model for toluene. We acknowledge that for the highly polarized charge-transfer (CT) states characteristic of TADF emitters, a state-specific solvation model (e.g., SS-PCM) is theoretically required for quantitative accuracy. However, such approaches are computationally prohibitive for HTS. The ALPB model is considered sufficient for screening as it has been shown that while the absolute solvation effect may be approximated, the relative ordering of molecules and the overall structure-property trends are conserved in the condensed phase.
II.2.3 Specific limitations for multi-resonance emitters
It is important to note that Multi-Resonance (MR) TADF emitters are inherently susceptible to multi-reference character [22], a feature that current single-reference based TD-DFT and its semi-empirical variants (sTDA-xTB/sTD-DFT-xTB) may not fully capture. Consequently, while our HTS trends provide excellent initial screening recommendations, definitive quantitative validation for MR-TADF candidates should rely on post-hoc applications of high-level methods such as SCS-CC2 or STEOM-DLPNO-CCSD [23].
II.3 Derivation of structure-property descriptors and theoretical framework
To translate the large computational dataset into actionable design rules, donor-acceptor (D-A) torsional angles were automatically extracted from all optimized geometries. To quantify the degree of charge-transfer and its relation to geometry, we calculated the HOMO-LUMO overlap integral ($S^{\prime}_{\text{HL}}$) and the distance between the centroids of the hole and electron densities ($\Delta r_{\text{CT}}$) using the Multiwfn package [24, 25]. Principal Component Analysis (PCA) was performed on the standardized matrix of all descriptors to identify the underlying variables governing TADF performance.
This analysis is grounded in the theoretical framework of TADF efficiency, which depends not only on a small $\Delta E_{\text{ST}}$(thermodynamics) but also on a high rate of reverse intersystem crossing ($k_{\text{RISC}}$), a kinetic parameter governed by spin-orbit coupling (SOC). According to El-Sayed’s rules, SOC is maximized when the interconverting singlet ($S_{1}$) and triplet ($T_{1}$) states have different orbital characters (e.g., charge-transfer vs. local-excitation) [26, 27]. Our chosen descriptors—particularly the D-A torsional angle—are direct proxies for tuning this CT/LE character balance. While our protocol does not compute SOC explicitly due to its prohibitive cost, analyzing these structural and electronic descriptors provides a robust, qualitative assessment of the kinetic factors that, alongside $\Delta E_{\text{ST}}$, govern overall TADF performance.
III Results and discussion: from statistical trends to rational design principles
This section leverages the large-scale dataset of $747$ TADF emitters, processed with the validated HTVS protocol detailed in our companion paper (Article 1)[13], to extract quantitative, statistically robust structure-property relationships. Our analysis transitions from broad statistical trends to specific, actionable molecular design rules, with a focus on providing a deep mechanistic understanding that connects molecular structure to the key photophysical parameters governing TADF efficiency.
III.1 Low-dimensional nature of the TADF design space
[…]
III.2 Architectural hierarchy defines TADF potential
[…]
Figure 2: TADF performance stratified by molecular architecture. The box plots show the distribution of the singlet-triplet energy gap ($\Delta E_{\text{ST}}$) for each class. D-A-D architectures demonstrate statistically superior characteristics (lower median and narrower distribution of $\Delta E_{\text{ST}}$) compared to other systems, confirming the efficacy of this design principle for consistently achieving small energy gaps.
III.3 Data-driven clustering reveals distinct molecular families
To complement the structure-based classification, we performed an unsupervised K-means clustering based on the molecules’ photophysical properties. This data-driven approach identifies four distinct molecular families, as shown in Figure 3:
[…]
  • Cluster 3 (<<1%, n=1): Outliers. A single outlier with extreme properties requiring individual investigation.
This analysis confirms that while structural classification is useful, the photophysical property space naturally segregates molecules into performance-based families, with a clear cluster of high-potential candidates emerging directly from the data.
Figure 3: K-means clustering (k=4) of the $747$ TADF emitters based on their standardized photophysical properties in both the gas phase (a) and toluene solvent (b). The analysis partitions the dataset into distinct families, most notably identifying a cluster of high-efficiency candidates (Cluster 1) with desirable properties for TADF applications.
III.4 Multi-resonance TADF: a distinct paradigm for blue emitters
[…]
However, as noted in the methodology, it is important to interpret these results with caution. The inherent electronic complexity of MR-TADF systems makes them prone to multi-reference character [22], which is not fully captured by the single-reference nature of the semi-empirical methods used in this HTS. The trends presented should therefore be interpreted as a valuable guide for compound ranking rather than a source of definitive quantitative predictions, with final validation requiring higher levels of theory.
Table 1: Photophysical property comparison of MR-TADF with conventional D-A systems, calculated with sTDA-xTB in toluene. MR emitters achieve a small $\Delta E_{\text{ST}}$comparable to D-A-D systems but with significantly higher oscillator strength and blue-shifted emission, highlighting a different design strategy.
Property MR-TADF (n=67) D-A-D (n=182) Simple D-A (n=248)
[…]
Oscillator Strength ($f$) $0.61$ p m 0.58 $0.40$ p m 0.52 $0.51$ p m 0.58
[…]
III.5 Deconstructing performance: geometric and electronic design levers
III.5.1 Torsional angle as a primary control parameter
[…]
III.5.2 Quantifying the torsion-induced electronic trade-off
This geometric control is directly reflected in the electronic structure. As shown in Figure 5, twisting the molecule reduces the HOMO-LUMO overlap ($S^{\prime}_{\text{HL}}$), which correlates strongly with a smaller $\Delta E_{\text{ST}}$(Spearman $\rho=-0.45$). However, this comes at a cost: the oscillator strength, essential for a high radiative rate, is directly proportional to this overlap ($\rho=0.78$). This visualization makes the fundamental TADF design trade-off explicit: minimizing the energy gap via orbital separation inherently suppresses emission intensity. An optimal region for high-performance emitters lies where there is intermediate overlap (0.15 to 0.300.150.30) that balances these competing requirements.
To move beyond this qualitative picture and establish a more rigorous quantitative link, future work could employ methods such as Natural Transition Orbitals (NTOs) or the RespA procedure [29, 30]. These tools would allow for a compact representation of the virtual excitations and rigorously connect the molecular geometry to the precise charge-transfer character (e.g., via centroid distance $\Delta r$) and the ensuing coupling efficiency [31].
Figure 5: Correlation analysis between HOMO-LUMO overlap and key photophysical properties. The strong negative correlation with $\Delta E_{\text{ST}}$(left) validates the spatial separation model for TADF design, while the strong positive correlation with oscillator strength (right) highlights the fundamental design trade-off that molecular engineers must navigate.
III.6 Implications for reverse intersystem crossing ($k_{\text{RISC}}$)
While a small $\Delta E_{\text{ST}}$is the primary thermodynamic requirement for TADF, the overall efficiency is governed by the kinetics of RISC. While direct calculation of the $k_{\text{RISC}}$ rate is outside the scope of HTS, the derived structural principles must be explicitly connected to the kinetics of TADF. Within the Marcus theory framework, $k_{\text{RISC}}$ is proportional to the square of the spin-orbit coupling (SOC) matrix element and exponentially dependent on $\Delta E_{\text{ST}}$and the reorganization energy ($\lambda$) [32].
According to El-Sayed’s rules, SOC is maximized when the interconverting singlet ($S_{1}$) and triplet ($T_{1}$) states have different orbital characters (e.g., charge-transfer vs. local-excitation, 1CT $\leftrightarrow$ 3LE) [26]. This creates a mechanistic challenge: the very act of creating a pure CT state in $S_{1}$ to minimize $\Delta E_{\text{ST}}$can lead to a $T_{1}$ state that also has pure CT character, making the transition spin-forbidden (1CT $\leftrightarrow$ 3CT) and suppressing $k_{\text{RISC}}$.
[…]
Table 2: Connection between molecular descriptors and the physical factors governing the RISC rate ($k_{\text{RISC}}$).
Molecular descriptor Influence on RISC factor Optimal range/strategy
HOMO-LUMO overlap Directly modulates $\Delta E_{\text{ST}}$ (thermodynamics) 0.15 to 0.300.150.30
[…]
Heteroatoms (S, Se) Directly increase SOC via heavy-atom effect Incorporate S-based donors
Conformational rigidity (e.g., in MR systems) Minimizes reorganization energy $\mathbf{\lambda}$ (kinetics) Use bridged or fused structures
III.7 Synthesis of a data-driven design strategy
The combination of these large-scale statistical analyses allows us to formulate a set of quantitative, data-driven design guidelines for accelerating the discovery of novel TADF emitters, summarized in Table 3. These rules provide a clear roadmap for molecular engineering, from selecting a high-performance architecture to fine-tuning its properties via geometric and electronic modifications.
Table 3: Quantitative TADF design guidelines derived from the high-throughput screening.
Parameter Target range / strategy Rationale
Architecture D-A-D or MR Statistically superior performance for general or blue emission, respectively.
[…]
HOMO-LUMO Overlap ($S^{\prime}_{\text{HL}}$) 0.15 to 0.300.150.30 Navigates the trade-off between $\Delta E_{\text{ST}}$  and oscillator strength.
Oscillator Strength ($f_{S_{1}}$) >0.1>0.1 Ensures a high radiative decay rate for bright emission.
III.8 Future outlook and data exploitation
III.8.1 The HTS dataset as a machine learning substrate
[…]
III.8.2 Actionable design targets
[…]
IV Conclusions
In this work, we performed a large-scale computational screening of $747$ TADF emitters to extract quantitative, statistically robust structure-property relationships for guiding molecular design. By systematically analyzing the impact of molecular architecture, conformational geometry, and electronic structure, we have moved beyond anecdotal evidence to establish a data-driven framework that integrates both the thermodynamic and kinetic principles of TADF engineering.
[…]
By providing a comprehensive set of design rules, a curated list of promising candidates, and a unified mechanistic view, this study significantly advances the goal of rational, predictive design for the next generation of TADF emitters. The experimental characterization of the identified candidates is the most important next step to validate our findings and realize the potential of this data-driven approach.
Conflicts of interest
There are no conflicts to declare.
V Associated content
V.1 Supporting information
The Supporting Information is available free of charge at https://pubs.acs.org/doi/...
Table 4: Molecular architecture and classification of the studied TADF emitters. This table provides the SMILES representation of each molecule, along with its structural classification.
Molecule Id SMILES Total donors Total acceptors Architecture Molecular weight logp
4CzTPN N#Cc1c(-n2c3ccccc3c3ccccc32)c(-n2c3ccccc3c3ccccc32)c(C#N)c(-n2c3ccccc3c3ccccc32)c1-n1c2ccccc2c2ccccc21 4 4 Multi-D/A $788.914$ $13.818$
DMAC-DPS CC1(C)c2ccccc2N(c2ccc(S(=O)(=O)c3ccc(N4c5ccccc5C(C)(C)c5ccccc54)cc3)cc2)c2ccccc21 2 0 2D-0A $632.829$ $10.738$
PSPCz C1=CC=C(C=C1)S(=O)(=O)C2=CC=CC(=C2)N3C4=CC=CC=C4C5=CC=CC=C53 1 0 1D-0A $383.472$ $5.617$
Px2BP C1=CC=C2C(=C1)N(C3=CC=CC=C3O2)C4=CC=C(C=C4)C(=O)C5=CC=C(C=C5)N6C7=CC=CC=C7OC8=CC=CC=C86 4 0 Multi-D/A $544.610$ $10.069$
CzS2 C1=CC=C2C(=C1)C3=CC=CC=C3N2C4=CC=C(C=C4)S(=O)(=O)C5=CC=C(C=C5)N6C7=CC=CC=C7C8=CC=CC=C86 2 0 2D-0A $548.667$ $8.714$
2TCz-DPS CC(C)(C)C1=CC2=C(C=C1)N(C3=C2C=C(C=C3)C(C)(C)C)C4=CC=C(C=C4)S(=O)(=O)C5=CC=C(C=C5)N6C7=C(C=C(C=C7)C(C)(C)C)C8=C6C=CC(=C8)C(C)(C)C 2 0 2D-0A $773.099$ $13.904$
TDBA-DI CC(C)(C)c1ccc2c(c1)B1c3cc(C(C)(C)C)ccc3Oc3cc(-n4c5ccccc5c5c6c(c7ccccc7n6-c6ccccc6)c6c(c7ccccc7n6-c6ccccc6)c54)cc(c31)O2 3 0 Multi-D/A $877.897$ $14.300$
OPDPO O=CC1=CCC(c2ccccc2N2c3ccccc3Sc3ccccc32)(P(=O)(c2ccccc2)c2ccccc2)C=C1 2 0 2D-0A $581.677$ $8.916$
2CzTPEPCz C(=C(c1ccc(-n2c3ccccc3c3ccccc32)cc1)c1ccc(-n2c3ccccc3c3ccccc32)cc1)c1ccc(-c2ccc(-n3c4ccccc4c4ccccc43)cc2)cc1 3 0 Multi-D/A $828.031$ $16.234$
TBRb CC(C)(C)c1ccc(-c2c3ccc(C(C)(C)C)cc3c(-c3ccccc3)c3c(-c4ccc(C(C)(C)C)cc4)c4ccc(C(C)(C)C)cc4c(-c4ccccc4)c23)cc1 0 0 Unknown $757.118$ $17.004$
TBPe CC(C)(C)c1cc2cc(C(C)(C)C)cc3c4cc(C(C)(C)C)cc5cc(C(C)(C)C)cc(c(c1)c23)c54 0 0 Unknown $476.748$ $10.927$
2CzPN N#Cc1cc(-c2cccc3c2[nH]c2ccccc23)c(-c2cccc3c2[nH]c2ccccc23)cc1C#N 2 4 Multi-D/A $458.524$ $8.033$
3CzFTFP Fc1nc(F)c(F)c(C2(c3ccc4c(c3)c3ccccc3n4-c3ccccc3)c3ccccc3-c3ccccc32)c1F 1 0 1D-0A $556.562$ $9.098$
DDPhCz-DCPP N#Cc1nc2c3ccc(-n4c5ccc(-c6ccccc6)cc5c5cc(-c6ccccc6)ccc54)cc3c3cc(-n4c5ccc(-c6ccccc6)cc5c5cc(-c6ccccc6)ccc54)ccc3c2nc1C#N 2 3 Multi-D/A $915.072$ $16.542$
ACRXTN CC1(C)c2ccccc2N(c2ccc3c(=O)c4ccccc4oc3c2)c2ccccc21 1 0 1D-0A $403.481$ $7.055$
TXO-TPA O=C1c2ccccc2S(=O)(=O)c2ccc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)cc21 1 0 1D-0A $487.580$ $7.201$
PCTXO O=C1c2ccccc2S(=O)(=O)c2ccc(-c3ccc(N(c4ccc(Br)cc4)c4ccc(Br)cc4)cc3)cc21 1 0 1D-0A $645.372$ $8.726$
IP-PPI c1ccc(-n2c(-c3ccc(-c4ccc(-c5cn6ccccc6n5)cc4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 0 0 Unknown $562.676$ $9.981$
DMDHNP-DPS CC1(C)c2cnccc2Nc2ccncc21.O=S(=O)(c1ccccc1)c1ccccc1 0 0 Unknown $429.545$ $5.379$
4NTAZ-PPI c1ccc(-c2nnc(-c3ccc(-c4ccc(-c5nc6c7ccccc7c7ccccc7c6n5-c5ccccc5)cc4)cc3)n2-c2cccc3ccccc23)cc1 0 0 Unknown $715.860$ $12.734$
TAZ-PPI c1ccc(-c2nnc(-c3ccc(-c4ccc(-c5nc6c7ccccc7c7ccccc7c6n5-c5ccccc5)cc4)cc3)n2-c2ccccc2)cc1 0 0 Unknown $665.800$ $11.581$
PTZ-mPYR Cc1cc(N2c3ccccc3Sc3ccccc32)ccc1-c1cc(-c2ccc(N3c4ccccc4Sc4ccccc43)cc2C)ncn1 4 1 Multi-D/A $654.864$ $12.296$
pCNBCzoCF3 N#Cc1ccc(-n2c3ccccc3c3cc(-c4ccc5c(c4)c4ccccc4n5-c4ccc(C#N)cc4C(F)(F)F)ccc32)c(C(F)(F)F)c1 2 4 Multi-D/A $670.616$ $11.329$
PTZTPE-2 CC1(C)OB(c2ccc(C(=C(c3ccccc3)c3ccccc3)c3ccccc3)cc2)OC1(C)C 0 0 Unknown $458.410$ $6.993$
4CzIPN N#Cc1c(-n2c3ccccc3c3ccccc32)c(C#N)c(-n2c3ccccc3c3ccccc32)c(-n2c3ccccc3c3ccccc32)c1-n1c2ccccc2c2ccccc21 4 4 Multi-D/A $788.914$ $13.818$
DMAC-TRZ CC1(C)c2ccccc2N(c2ccc(-c3nc(-c4ccccc4)nc(-c4ccccc4)n3)cc2)c2ccccc21 1 1 D-A $516.648$ $8.982$
TCzTrz c1ccc(-c2nc(-c3ccccc3)nc(-c3cc(-n4c5ccccc5c5ccccc54)c(-n4c5ccccc5c5ccccc54)c(-n4c5ccccc5c5ccccc54)c3)n2)cc1 3 1 Multi-D/A $804.957$ $14.164$
$\delta$-2CbPN N#Cc1cc(-n2c3ccccc3c3ncccc32)c(-n2c3ccccc3c3ncccc32)cc1C#N 0 4 Multi-D/A $460.500$ $6.414$
CzDCbTrz c1ccc(-c2nc(-c3ccccc3)nc(-c3cc(-n4c5ccccc5c5ncccc54)c(-n4c5ccccc5c5ccccc54)c(-n4c5ccccc5c5ncccc54)c3)n2)cc1 1 1 D-A $806.933$ $12.954$
$\alpha$-2CbPN N#Cc1cc(-n2c3ccccc3c3cccnc32)c(-n2c3ccccc3c3cccnc32)cc1C#N 0 4 Multi-D/A $460.500$ $6.414$
CN-QP N#Cc1ccc2nc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)nc2c1 4 3 Multi-D/A $669.744$ $11.986$
TFM-QP FC(F)(F)c1ccc2nc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)nc2c1 4 1 Multi-D/A $712.731$ $13.134$
NDN c1ccc2c(c1)c1cccnc1n2-c1ccc2sc3ccc(-n4c5ccccc5c5cccnc54)cc3c2c1 0 0 Unknown $516.629$ $9.039$
mCP c1cc(-n2c3ccccc3c3ccccc32)cc(-n2c3ccccc3c3ccccc32)c1 2 0 2D-0A $408.504$ $7.881$
ZDZ c1ccc2c(c1)c1ccccc1n2-c1ccc2sc3ccc(-n4c5ccccc5c5ccccc54)cc3c2c1 2 0 2D-0A $514.653$ $10.249$
ZDN c1ccc2c(c1)c1ccccc1n2-c1ccc2sc3ccc(-n4c5ccccc5c5cccnc54)cc3c2c1 1 0 1D-0A $515.641$ $9.644$
BTCz3CzTPN N#Cc1c(-n2c3ccccc3c3ccccc32)c(-n2c3ccccc3c3ccccc32)c(C#N)c(-n2c3ccccc3c3c4sc5ccccc5c4ccc32)c1-n1c2ccccc2c2ccccc21 4 4 Multi-D/A $895.063$ $16.186$
2BTCz2CzTPN N#Cc1c(-n2c3ccccc3c3ccccc32)c(-n2c3ccccc3c3c4sc5ccccc5c4ccc32)c(C#N)c(-n2c3ccccc3c3ccccc32)c1-n1c2ccccc2c2c3sc4ccccc4c3ccc21 4 4 Multi-D/A 1 001.2121\,001.212 $18.554$
APDC-DTPA N#Cc1nc2c(nc1C#N)-c1ccc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)c3c(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)ccc-2c13 2 2 2D-2A $740.870$ $13.294$
BMIM N#Cc1nc2c(nc1C#N)-c1ccc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)c3c(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)ccc-2c13 2 2 2D-2A $740.870$ $13.294$
ADO-TPA O=C1C(=O)c2ccc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)c3cccc1c23 1 0 1D-0A $425.487$ $7.356$
DCzTrz c1ccc(-c2nc(-c3ccccc3)nc(-c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)n2)cc1 2 1 D-A-D $639.762$ $11.067$
BDPCC c1ccc(-c2ccc3[nH]c4c(-c5ccc6c([nH]c7ccccc76)c5-c5cc(-c6ccccc6)cc6c5[nH]c5ccc(-c7ccccc7)cc56)cc(-c5ccccc5)cc4c3c2)cc1 3 0 Multi-D/A $801.993$ $16.592$
SpiroAC-TRZ c1ccc(-c2nc(-c3ccccc3)nc(-c3ccc(N4c5ccccc5C5(c6ccccc6-c6ccccc65)c5ccccc54)cc3)n2)cc1 1 1 D-A $638.774$ $11.019$
DBT-BZ-DMAC CC1(C)c2ccccc2Nc2cccc(C(=O)c3ccccc3)c21.c1ccc2c(c1)sc1ccccc12 0 0 Unknown $497.663$ $9.355$
DQP c1ccc2nc3c(nc2c1)c1nc2ccccc2nc1c1nc2ccccc2nc31 0 6 Multi-D/A $384.402$ $4.976$
TPA-DCPP N#Cc1nc2c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cc3c3cc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)ccc3c2nc1C#N 2 3 Multi-D/A $766.908$ $13.953$
2PXZ-OXD c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2nnc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)o2)cc1 4 0 Multi-D/A $584.635$ $10.555$
BmPyPhB CC1(C)c2ccccc2N(c2ccc(-c3nc(-c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)nc(-c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)n3)cc2)c2ccccc21 3 1 Multi-D/A $931.200$ $17.200$
MPPA COc1ccc(N(c2ccc(OC)cc2)c2ccc(-c3ccc4c(c3)C(=O)c3ccc(-c5ccc(N(c6ccc(OC)cc6)c6ccc(OC)cc6)cc5)cc3C4=O)cc2)cc1 2 0 2D-0A $814.938$ $12.770$
ACRFLCN N#Cc1ccc2c(c1)C1(c3cc(C#N)ccc3-2)c2ccccc2N(c2ccccc2)c2ccccc21 1 4 Multi-D/A $457.536$ $7.576$
PTZ-DCPP N#Cc1nc2c3ccc(N4c5ccccc5Sc5ccccc54)cc3c3cc(N4c5ccccc5Sc5ccccc54)ccc3c2nc1C#N 4 3 Multi-D/A $674.814$ $11.548$
BBA-Pyr c1ccc(-c2ccc(N(c3ccc(-c4ccccc4)cc3)c3ccc(-c4ccc(-c5cc(-c6ccccc6)nc(-c6ccccc6)n5)cc4)cc3)cc2)cc1 1 1 D-A $703.889$ $13.948$
PhICz-Pyr c1ccc(-c2ccc3c(c2)c2cc(-c4ccccc4)cc4c5cc(-c6ccc(-c7cc(-c8ccccc8)nc(-c8ccccc8)n7)cc6)ccc5n3c24)cc1 2 1 D-A-D $699.857$ $13.629$
PhCz-Pyr c1ccc(-c2ccc3c(c2)c2cc(-c4ccccc4)ccc2n3-c2ccc(-c3ccc(-c4cc(-c5ccccc5)nc(-c5ccccc5)n4)cc3)cc2)cc1 1 1 D-A $701.873$ $13.576$
PDA-DP CC1(C)c2cc(-c3ccccc3)ccc2N(c2ccc(S(=O)(=O)c3ccc(N4c5ccc(-c6ccccc6)cc5C(C)(C)c5cc(-c6ccccc6)ccc54)cc3)cc2)c2ccc(-c3ccccc3)cc21 2 0 2D-0A $937.221$ $17.406$
9CzFDBFDPO O=P(c1ccccc1)(c1ccccc1)c1cccc2c1oc1c(P(=O)(c3ccccc3)c3ccccc3)cc(-c3ccc(C4(c5ccc(-n6c7ccccc7c7ccccc76)cc5)c5ccccc5-c5ccccc54)cc3)cc12 1 0 1D-0A 1 050.1471\,050.147 $15.992$
9PhCzFDBFDPO O=P(c1ccccc1)(c1ccccc1)c1cccc2c1oc1c(P(=O)(c3ccccc3)c3ccccc3)cc(-c3ccc(C4(c5ccc(-c6ccc(-n7c8ccccc8c8ccccc87)cc6)cc5)c5ccccc5-c5ccccc54)cc3)cc12 1 0 1D-0A 1 126.2451\,126.245 $17.659$
9PhCzFDBFSPO O=P(c1ccccc1)(c1ccccc1)c1cccc2c1oc1ccc(-c3ccc(C4(c5ccc(-c6ccc(-n7c8ccccc8c8ccccc87)cc6)cc5)c5ccccc5-c5ccccc54)cc3)cc12 1 0 1D-0A $926.068$ $16.020$
9CzFDBFSPO O=P(c1ccccc1)(c1ccccc1)c1cccc2c1oc1ccc(-c3ccc(C4(c5ccc(-n6c7ccccc7c7ccccc76)cc5)c5ccccc5-c5ccccc54)cc3)cc12 1 0 1D-0A $849.970$ $14.353$
9ArFDBFxPO O=P(c1ccccc1)(c1ccccc1)c1cccc2c1oc1c(P(=O)(c3ccccc3)c3ccccc3)cc(-c3ccc(C4(c5ccc(-n6c7ccccc7c7ccccc76)cc5)c5ccccc5-c5ccccc54)cc3)cc12 1 0 1D-0A 1 050.1471\,050.147 $15.992$
12BTCzTPN N#Cc1c(-n2c3ccccc3c3ccccc32)c(-n2c3ccccc3c3ccc4c5ccccc5sc4c32)c(C#N)c(-n2c3ccccc3c3ccccc32)c1-n1c2ccccc2c2ccc3c4ccccc4sc3c21 4 4 Multi-D/A 1 001.2121\,001.212 $18.554$
oDTBPZ-DPXZ Cc1ccccc1-c1cc2nc3c4ccc(N5c6ccccc6Oc6ccccc65)cc4c4cc(N5c6ccccc6Oc6ccccc65)ccc4c3nc2cc1-c1ccccc1C 4 2 Multi-D/A $822.968$ $16.191$
3DMAC-BP-CN CC1(C)c2ccccc2N(c2ccc3c(c2)c2cc(N4c5ccccc5C(C)(C)c5ccccc54)ccc2c2nc4cc(N5c6ccccc6C(C)(C)c6ccccc65)c(C#N)cc4nc32)c2ccccc21 3 4 Multi-D/A $927.168$ $17.288$
3DMAC-BP-Br CC1(C)c2ccccc2N(c2ccc3c(c2)c2cc(N4c5ccccc5C(C)(C)c5ccccc54)ccc2c2nc4cc(N5c6ccccc6C(C)(C)c6ccccc65)c(Br)cc4nc32)c2ccccc21 3 2 Multi-D/A $981.054$ $18.179$
TPA-PZCN N#Cc1ccc2c(c1)c1cc(C#N)ccc1c1nc3cc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)c(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cc3nc21 2 6 Multi-D/A $816.968$ $15.106$
1PXZ-BP c1ccc2c(c1)Oc1ccccc1N2c1ccc2c(c1)c1ccccc1c1nc3ccccc3nc21 2 2 2D-2A $461.524$ $8.665$
PTZ-DBTO2 O=S1(=O)c2ccccc2-c2cc(N3c4ccccc4Sc4ccccc43)ccc21 2 0 2D-0A $413.523$ $6.434$
BACQ CC1(C)c2c(ccc3ccccc23)-n2c3ccccc3c(=O)c3cccc1c32 1 0 1D-0A $361.444$ $5.936$
2Cz-DMAC-TXO CC1(C)c2cc(-n3c4ccccc4c4ccccc43)ccc2N(c2ccc3sc4ccccc4c(=O)c3c2)c2ccc(-n3c4ccccc4c4ccccc43)cc21 3 0 Multi-D/A $749.939$ $13.718$
2Cz-DMAC-TTR CC1(C)c2cc(-n3c4ccccc4c4ccccc43)ccc2N(c2ccc3c(c2)S(=O)(=O)c2ccccc2S3(=O)=O)c2ccc(-n3c4ccccc4c4ccccc43)cc21 3 0 Multi-D/A $817.992$ $11.969$
2Cz-DMAC-BTB CC1(C)c2cc(-n3c4ccccc4c4ccccc43)ccc2N(c2ccc3c(c2)-c2ccccc2S3(=O)=O)c2ccc(-n3c4ccccc4c4ccccc43)cc21 3 0 Multi-D/A $753.927$ $12.803$
DMAC-BPP CC1(C)c2ccccc2N(c2cc(-c3cc(-c4ccc(-c5ccccc5)cc4)nc(-c4ccccc4)n3)cc(N3c4ccccc4C(C)(C)c4ccccc43)c2)c2ccccc21 2 1 D-A-D $799.034$ $15.363$
Me-MOC COc1ccc2c(c1)c1ccccc1n2-c1ccc(-c2c(C#N)c(C)nc(C)c2C#N)cc1 1 2 A-D-A $428.495$ $6.215$
Me-DOC COc1ccc2c(c1)c1cc(OC)ccc1n2-c1ccc(-c2c(C#N)c(C)nc(C)c2C#N)cc1 1 2 A-D-A $458.521$ $6.223$
3DCPO O=P(c1ccccc1)(c1ccc2c(c1)c1ccccc1n2-c1ccccc1)c1ccc2c(c1)c1ccccc1n2-c1ccccc1 2 0 2D-0A $608.681$ $9.520$
TDBA-TPDICz CC(C)(C)c1ccc2c(c1)B1c3cc(C(C)(C)C)ccc3Oc3cc(-n4c5ccccc5c5c(-c6ccccc6)c6c(c(-c7ccccc7)c54)c4ccccc4n6-c4ccccc4)cc(c31)O2 2 0 2D-0A $864.898$ $14.537$
TRZ-pIC CC(C)(C)c1ccc2c(c1)B1c3cc(C(C)(C)C)ccc3Oc3cc(-n4c5ccccc5c5c(-c6ccccc6)c6c(c(-c7ccccc7)c54)c4ccccc4n6-c4ccccc4)cc(c31)O2 2 0 2D-0A $864.898$ $14.537$
SBF-PXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2c(c1)C1(c3ccccc3-c3ccccc31)c1cc(N3c4ccccc4Oc4ccccc43)ccc1-2 4 0 Multi-D/A $678.791$ $13.181$
PtOEP CCC1=C($C$)c2cc3[nH]c(cc4nc(cc5[nH]c(cc1n2)c($C$)c5CC)C(CC)=C4CC)c($C$)c3CC.[Pt+2] 0 0 Unknown $729.870$ $10.023$
PxCYP N#Cc1ccc(-c2ccc(N3c4ccccc4Oc4ccccc43)cc2)nc1-c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 1 Multi-D/A $618.696$ $11.438$
AcCYP CC1(C)c2ccccc2N(c2ccc(-c3ccc(C#N)c(-c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)n3)cc2)c2ccccc21 2 1 D-A-D $670.860$ $12.505$
BCN CC1(C)c2ccccc2N2c3ccc4cc(C#N)ccc4c3C(C)(C)c3cccc1c32 1 2 A-D-A $400.525$ $7.460$
DPPSPF O=S(=O)(c1ccccc1)c1ccc(-c2ccc3c(c2)C(c2ccccc2)(c2ccccc2)c2ccccc2-3)cc1 0 0 Unknown $534.680$ $8.550$
Me2ACBO CC(C)(c1ccc2c(c1)B1c3ccc(N4c5ccccc5C(C)(C)c5ccccc54)cc3Oc3cccc(c31)O2)c1ccc2c(c1)B1c3ccc(N4c5ccccc5C(C)(C)c5ccccc54)cc3Oc3cccc(c31)O2 2 0 2D-0A $994.809$ $13.686$
F2AcBO CC1(C)c2ccccc2N(c2ccc3c(c2)Oc2cccc4c2B3c2cc(C3(c5ccc6c(c5)B5c7ccc(N8c9ccccc9C(C)(C)c9ccccc98)cc7Oc7cccc(c75)O6)c5ccccc5-c5ccccc53)ccc2O4)c2ccccc21 2 0 2D-0A 1 116.9351\,116.935 $15.723$
PSPSF O=S(=O)(c1ccccc1)c1ccc(-c2ccc3c(c2)C2(c4ccccc4-c4ccccc42)c2ccccc2-3)cc1 0 0 Unknown $532.664$ $8.530$
TXO-P-Si O=C1c2ccccc2S(=O)(=O)c2ccc(-c3ccc([Si](c4ccccc4)(c4ccccc4)c4ccccc4)cc3)cc21 0 0 Unknown $578.765$ $5.108$
CzB-FMPIM FC(F)(F)c1ccc(-n2c(-c3ccc(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)cc3)nc(-c3ccccc3)c2-c2ccccc2)cc1 1 0 1D-0A $681.761$ $12.656$
CzB-FMPPI FC(F)(F)c1ccc(-n2c(-c3ccc(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 1 0 1D-0A $679.745$ $12.782$
BACF CC1(C)c2ccccc2N(c2ccccc2)c2ccc3cc(F)ccc3c21 1 0 1D-0A $353.440$ $7.088$
BACH CC1(C)c2ccccc2N(c2ccccc2)c2ccc3ccccc3c21 1 0 1D-0A $335.450$ $6.949$
PzTDBA CC(C)(C)c1ccc2c(c1)B1c3cc(C(C)(C)C)ccc3Oc3cc(N4c5ccccc5N(c5cc6c7c(c5)Oc5ccc(C(C)(C)C)cc5B7c5cc(C(C)(C)C)ccc5O6)c5ccccc54)cc(c31)O2 2 0 2D-0A $942.818$ $13.585$
3TCPM O=C(c1ccccc1)c1ccc(-n2c3ccccc3c3cc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)ccc32)cc1 2 0 2D-0A $590.726$ $11.151$
TPA-PPO c1ccc(-c2nc(-c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cn3)oc2-c2ccccc2)cc1 1 0 1D-0A $541.654$ $10.207$
CzPh-PPO c1ccc(-c2nc(-c3ccc(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)cn3)oc2-c2ccccc2)cc1 1 0 1D-0A $539.638$ $9.835$
m-CF3PIBI FC(F)(F)c1cccc(-n2c(-c3ccc(-c4ccc(-c5ccc(-c6nc7ccccc7n6-c6ccccc6)cc5)cc4)cc3)nc3c4ccccc4c4ccccc4c32)c1 0 0 Unknown $782.869$ $14.358$
DPTPEA c1ccc(C(=C(c2ccccc2)c2ccc(N(c3ccccc3)c3ccccc3)cc2)c2ccccc2)cc1 1 0 1D-0A $499.657$ $10.164$
PICNAnCz N#Cc1ccc(-n2c(-c3ccc(-c4c5ccccc5c(Br)c5ccccc45)cc3)nc3c4ccccc4c4ccccc4c32)cc1 0 2 0D-2A $650.579$ $11.606$
DPA-DPQ c1ccc(-c2cc3ncc(N(c4ccccc4)c4ccccc4)nc3cc2-c2ccccc2)cc1 0 1 0D-1A $449.557$ $8.434$
PhQLPXZ c1ccc(-c2cc(-c3ccccc3)c3cc(N4c5ccccc5Oc5ccccc54)ccc3n2)cc1 2 0 2D-0A $462.552$ $9.144$
2DPA-DPQ c1ccc(-c2cc3nc(N(c4ccccc4)c4ccccc4)c(N(c4ccccc4)c4ccccc4)nc3cc2-c2ccccc2)cc1 0 1 0D-1A $616.768$ $11.903$
FDQPXZ Fc1cc2nc(-c3ccccc3)c(-c3ccccc3)nc2cc1N1c2ccccc2Oc2ccccc21 2 1 D-A-D $481.530$ $8.678$
DTRZSFX c1ccc(-c2nc(-c3ccccc3)nc(-c3ccc4c(c3)C3(c5ccccc5Oc5ccccc53)c3cc(-c5nc(-c6ccccc6)nc(-c6ccccc6)n5)ccc3-4)n2)cc1 0 2 0D-2A $794.918$ $12.527$
E-o-DHMPA OCc1ccccc1-c1c2ccccc2c(-c2ccccc2CO)c2ccccc12 0 0 Unknown $390.482$ $6.312$
Z-o-DHMPA OCc1ccccc1-c1c2ccccc2c(-c2ccccc2CO)c2ccccc12 0 0 Unknown $390.482$ $6.312$
p-DHMPA OCc1ccc(-c2c3ccccc3c(-c3ccc(CO)cc3)c3ccccc23)cc1 0 0 Unknown $390.482$ $6.312$
STAC-BP-STAC O=C(c1ccc(N2c3ccccc3C3(c4ccccc4Sc4ccccc43)c3ccccc32)cc1)c1ccc(N2c3ccccc3C3(c4ccccc4Sc4ccccc43)c3ccccc32)cc1 2 0 2D-0A $905.160$ $16.179$
SXAC-BP-SXAC O=C(c1ccc(N2c3ccccc3C3(c4ccccc4Oc4ccccc43)c3ccccc32)cc1)c1ccc(N2c3ccccc3C3(c4ccccc4Oc4ccccc43)c3ccccc32)cc1 2 0 2D-0A $873.024$ $15.461$
m-DHMPA OCc1cccc(-c2c3ccccc3c(-c3cccc(CO)c3)c3ccccc23)c1 0 0 Unknown $390.482$ $6.312$
SFAC-BP-SFAC O=C(c1ccc(N2c3ccccc3C3(c4ccccc4-c4ccccc43)c3ccccc32)cc1)c1ccc(N2c3ccccc3C3(c4ccccc4-c4ccccc43)c3ccccc32)cc1 2 0 2D-0A $841.026$ $15.210$
PBTPA CCN1c2ccccc2Sc2cc(-c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)c4nsnc34)ccc21 2 0 2D-0A $604.804$ $11.118$
HDP O=C1c2ccccc2C(=O)c2nc3c(cc(O)c4ccccc43)nc21 0 1 0D-1A $326.311$ $3.264$
PY-TPA CC(C)(C)c1ccc(N(c2ccc(-c3cc4nc5c6ccc(C#N)cc6c6cc(C#N)ccc6c5nc4cc3-c3cccnc3)cc2)c2ccc(C(C)(C)C)cc2)cc1 1 6 Multi-D/A $762.961$ $13.627$
CN-TPA CC(C)(C)c1ccc(N(c2ccc(-c3cc4nc5c6ccc(C#N)cc6c6cc(C#N)ccc6c5nc4cc3C#N)cc2)c2ccc(C(C)(C)C)cc2)cc1 1 8 Multi-D/A $710.885$ $12.436$
DMIC-TRZ CC1(C)CC(c2cccc(-c3nc(-c4ccccc4)nc(-c4ccccc4)n3)c2)C=c2[nH]c3cc4c(cc3c21)-c1ccccc1C=4 0 1 0D-1A $592.746$ $8.409$
NyDPAc CC1(C)c2ccccc2Nc2cccc(-c3ccc(-c4ccc5nc(-c6ccc(-c7cccc8c7C(C)(C)c7ccccc7N8)cc6)ccc5n4)cc3)c21 0 0 Unknown $696.898$ $13.064$
oCBP c1ccc(-n2c3ccccc3c3ccccc32)c(-c2ccccc2-n2c3ccccc3c3ccccc32)c1 2 0 2D-0A $484.602$ $9.548$
PABP Cc1cc(-c2c3ccccc3c(-c3ccccc3)c3ccccc23)ccc1-c1ccc(-c2c3ccccc3c(-c3ccccc3)c3ccccc23)cc1C 0 0 Unknown $686.898$ $15.251$
DHID-DPS CCn1c2ccccc2c2c1c1ccccc1n2-c1ccc(S(=O)(=O)c2ccc(-n3c4ccccc4c4c3c3ccccc3n4CC)cc2)cc1 0 0 Unknown $682.849$ $10.663$
4FBTPI CC(C)(C)c1ccc(-n2c(-c3ccc(-c4c($F)c($-c5ccc(-c6nc7c8ccccc8c8ccccc8c7n6-c6ccc(C(C)(C)C)cc6)cc5)c(F)c4F)cc3)nc3c4ccccc4c4ccccc4c32)cc1 0 0 Unknown $999.168$ $18.797$
BCzBMes Cc1cc(C)c(B(c2c(C)cc(C)cc2C)n2c3ccccc3c3cc(-c4ccc5c(c4)c4ccccc4n5-c4ccccc4)ccc32)c(C)c1 2 0 2D-0A $656.682$ $11.063$
CCDMB Cc1cc(C)c(B(c2ccc(-n3c4ccccc4c4cc(-n5c6ccccc6c6ccccc65)ccc43)cc2)c2c(C)cc(C)cc2C)c(C)c1 2 0 2D-0A $656.682$ $10.248$
PCDMB Cc1cc(C)c(B(c2ccc(-n3c4ccccc4c4cc(N5c6ccccc6Oc6ccccc65)ccc43)cc2)c2c(C)cc(C)cc2C)c(C)c1 3 0 Multi-D/A $672.681$ $10.726$
BTrztCz CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccccc1-c1nc(-c2ccccc2)nc(-c2ccc(-c3nc(-c4ccccc4)nc(-c4ccccc4-n4c5ccc(C(C)(C)C)cc5c5cc(C(C)(C)C)ccc54)n3)cc2)n1 2 2 2D-2A 1 095.4521\,095.452 $19.443$
DHID-DBS CCCCn1c2ccccc2c2c1c1ccccc1n2-c1ccc2c(c1)-c1cc(-n3c4ccccc4c4c3c3ccccc3n4CCCC)ccc1S2(=O)=O 0 0 Unknown $736.941$ $12.204$
CZ-DPS-DMAC CC1(C)c2ccccc2N(c2ccc(S(=O)(=O)c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)cc2)c2ccccc21 3 0 Multi-D/A $755.943$ $12.823$
XT-T Cc1ccc(N(c2ccc(C)cc2)c2ccccc2C2(c3ccccc3N(c3ccc(C)cc3)c3ccc(C)cc3)CCCCC2)cc1 2 0 2D-0A $626.888$ $13.110$
3DPTIA CC1(C)c2ccccc2N(c2nc(-c3ccccc3)nc(-c3ccccc3)n2)c2cc3c4ccccc4n(-c4ccccc4)c3cc21 1 1 D-A $605.745$ $10.412$
PyIAnTPh c1ccc(-c2cc(-c3ccccc3)cc(-c3c4ccccc4c(-c4ccc(-c5nc6c7cccc8ccc9cccc(c9c87)c6n5-c5ccccc5)cc4)c4ccccc34)c2)cc1 0 0 Unknown $798.989$ $16.564$
2CzP-INN N#Cc1c(-c2ccc(-n3c4ccccc4c4ccccc43)cc2)cncc1-c1ccc(-n2c3ccccc3c3ccccc32)cc1 2 1 D-A-D $586.698$ $10.481$
VBN0 C=Cc1ccc(COc2ccc3c(c2)c2ccccc2n3-c2cccc(-n3c4ccccc4c4cc(OCc5ccc(C=C)cc5)ccc43)c2C#N)cc1 2 2 2D-2A $697.838$ $12.196$
PIAnTPh c1ccc(-c2cc(-c3ccccc3)cc(-c3c4ccccc4c(-c4ccc(-c5nc6c7ccccc7c7ccccc7c6n5-c5ccccc5)cc4)c4ccccc34)c2)cc1 0 0 Unknown $774.967$ $15.973$
BCZ-DPS-AD CC1(C)c2ccccc2N(c2cc(N3c4ccccc4C(C)(C)c4ccccc43)cc(S(=O)(=O)c3ccc4c(c3)c3ccccc3n4-c3ccccc3)c2)c2ccccc21 3 0 Multi-D/A $798.024$ $13.835$
2BTPA-INN N#Cc1c(-c2ccc(N(c3ccc(-c4ccccc4)cc3)c3ccc(-c4ccccc4)cc3)cc2)cncc1-c1ccc(N(c2ccc(-c3ccccc3)cc2)c2ccc(-c3ccccc3)cc2)cc1 2 1 D-A-D $895.122$ $17.895$
SpiroAc c1ccc2c(c1)Nc1ccccc1C21c2ccccc2-c2ccccc21 0 0 Unknown $331.418$ $6.107$
BuPCzBCO Cc1ccc2c3cc(Br)c(C)cc3n(-c3ccc(C(C)(C)C)cc3)c2c1 1 0 1D-0A $406.367$ $7.461$
AQ-S-CZ O=C1c2ccccc2C(=O)c2cc(Sc3ccc(-n4c5ccccc5c5ccccc54)cc3)ccc21 1 0 1D-0A $481.576$ $7.710$
DBP-3TPA c1ccc(N(c2ccccc2)c2ccc(-c3ccc4nc5c6ccc(-c7ccc(N(c8ccccc8)c8ccccc8)cc7)cc6c6cc(-c7ccc(N(c8ccccc8)c8ccccc8)cc7)ccc6c5nc4c3)cc2)cc1 3 2 Multi-D/A 1 010.2571\,010.257 $20.500$
DCzIPN N#Cc1cc(C#N)c(-n2c3ccccc3c3ccccc32)cc1-n1c2ccccc2c2ccccc21 2 4 Multi-D/A $458.524$ $7.624$
DMeCzIPN Cc1cccc2c3ccccc3n(-c3cc(-n4c5ccccc5c5cccc(C)c54)c(C#N)cc3C#N)c12 2 4 Multi-D/A $486.578$ $8.241$
SAFDPA c1ccc(N(c2ccccc2)c2ccc3c(c2)-c2ccccc2C32c3ccccc3N(c3ccccc3)c3ccccc32)cc1 2 0 2D-0A $574.727$ $11.303$
t-DABNA CC(C)(C)c1ccc(N2c3ccc(C(C)(C)C)cc3B3c4cc(C(C)(C)C)ccc4N(c4ccc(C(C)(C)C)cc4)c4cccc2c43)cc1 2 0 2D-0A $644.756$ $10.959$
3-AcBPtCN CC1(C)c2ccccc2N(c2ccc3c(c2)c2cc(C#N)ccc2c2nc4cc(C#N)c(C#N)cc4nc32)c2ccccc21 1 8 Multi-D/A $562.636$ $8.814$
3-AcBPdCN CC1(C)c2ccccc2N(c2ccc3c(c2)c2ccccc2c2nc4cc(C#N)c(C#N)cc4nc32)c2ccccc21 1 6 Multi-D/A $537.626$ $8.942$
11-AcBPdCN CC1(C)c2ccccc2N(c2ccc3nc4c5ccc(C#N)cc5c5cc(C#N)ccc5c4nc3c2)c2ccccc21 1 6 Multi-D/A $537.626$ $8.942$
PID-3 c1ccc(-n2c(-c3ccc(-c4cccs4)s3)nc3c4ccccc4c4ccccc4c32)cc1 0 0 Unknown $458.611$ $8.789$
3-ACBPmCN CC1(C)c2ccccc2N(c2ccc3c(c2)c2cc(C#N)ccc2c2nc4ccccc4nc32)c2ccccc21 1 4 Multi-D/A $512.616$ $9.070$
3-DTPA-BBI O=c1c2ccc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)c3c(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)ccc(c32)c2nc3ccccc3n12 2 0 2D-0A $756.909$ $13.865$
IABTCN CC1(C)c2ccccc2-c2cc3c(cc21)N(c1ccccc1)c1ccc(-c2ccc(C#N)c4nsnc24)cc1C3(c1ccccc1)c1ccccc1 1 2 A-D-A $684.868$ $11.702$
BTZPP CC(C)(C)c1ccc(-n2c(-c3ccc(-c4cccc5nsnc45)cc3)cc3c2cc(-c2ccc(-c4cccc5nsnc45)cc2)n3-c2ccc(C(C)(C)C)cc2)cc1 0 0 Unknown $791.062$ $13.694$
[…]
BCzPh c1ccc(-n2c3ccccc3c3cc(-c4ccc5c(c4)c4ccccc4n5-c4ccccc4)ccc32)cc1 2 0 2D-0A $484.602$ $9.548$
TBPM2CN N#Cc1ccc(-c2ccc3c(c2)c2cc(-c4ccc(C#N)cc4)ccc2c2c3nc(-c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cc3)n2-c2ccccc2)cc1 1 4 Multi-D/A $815.980$ $15.213$
TPM2CN N#Cc1ccc(-c2ccc3c(c2)c2cc(-c4ccc(C#N)cc4)ccc2c2c3nc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)n2-c2ccccc2)cc1 1 4 Multi-D/A $739.882$ $13.546$
PT-BZ-DMAC CC1(C)c2ccccc2N(c2ccc(C(=O)c3ccc(Sc4ccccc4)cc3)cc2)c2ccccc21 1 0 1D-0A $497.663$ $9.178$
PS-BZ-DMAC CC1(C)c2ccccc2N(c2ccc(C(=O)c3ccc(S(=O)(=O)c4ccccc4)cc3)cc2)c2ccccc21 1 0 1D-0A $529.661$ $7.860$
2PCzBN-tPh CC(C)(C)c1ccc(-c2cc(-n3c4ccc(-c5ccccc5)cc4c4cc(-c5ccccc5)ccc43)c(C#N)c(-n3c4ccc(-c5ccccc5)cc4c4cc(-c5ccccc5)ccc43)c2)cc1 2 2 2D-2A $870.112$ $17.385$
2PCzBN N#Cc1c(-n2c3ccc(-c4ccccc4)cc3c3cc(-c4ccccc4)ccc32)cccc1-n1c2ccc(-c3ccccc3)cc2c2cc(-c3ccccc3)ccc21 2 2 2D-2A $737.906$ $14.420$
TPA-2 O=C1c2ccccc2C(=O)c2c1ccc1nc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)c(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)nc21 2 1 D-A-D $746.870$ $12.679$
36DTPAFM N#CC(C#N)=C1c2ccc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)cc2-c2cc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)ccc21 2 2 2D-2A $714.872$ $13.790$
AQTC N#Cc1cc2nc3c(nc2c(C#N)c1C#N)-c1cccc2cccc-3c12 0 7 Multi-D/A $329.322$ $4.045$
3TPAFM N#CC(C#N)=C1c2ccccc2-c2cc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)ccc21 1 2 A-D-A $471.563$ $8.653$
BBPA CCCCOc1ccccc1-c1c2ccccc2c(-c2ccccc2OCCCC)c2ccccc12 0 0 Unknown $474.644$ $9.685$
DMPA COc1ccc(-c2c3ccccc3c(-c3ccc(OC)cc3OC)c3ccccc23)c(OC)c1 0 0 Unknown $450.534$ $7.361$
PXZ-NAI O=C1c2cccc3c(N4c5ccccc5Oc5ccccc54)ccc(c23)C(=O)N1c1ccccc1 2 0 2D-0A $454.485$ $7.216$
PTZ-NAI O=C1c2cccc3c(N4c5ccccc5Sc5ccccc54)ccc(c23)C(=O)N1c1ccccc1 2 0 2D-0A $470.553$ $7.575$
AI-4Cz O=C1c2c(c(-n3c4ccccc4c4ccccc43)c(-n3c4ccccc4c4ccccc43)c(-n3c4ccccc4c4ccccc43)c2-n2c3ccccc3c3ccccc32)C(=O)N1c1ccccc1 4 0 Multi-D/A $884.011$ $14.876$
DAcIPN CC1(C)c2ccccc2N(c2cc(N3c4ccccc4C(C)(C)c4ccccc43)c(C#N)cc2C#N)c2ccccc21 2 4 Multi-D/A $542.686$ $9.648$
TCz1 CCCCC($C$)Cn1c2ccc(-n3c4ccccc4c4ccccc43)cc2c2cc(-n3c4ccccc4c4ccccc43)ccc21 3 0 Multi-D/A $609.817$ $12.205$
MBF F[B-]1(F)OC(c2ccccc2)=Cc2cc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)nc[n+]21 2 1 D-A-D $503.317$ $7.363$
MPA O=C(Cc1cc(-c2ccc(N3c4ccccc4Oc4ccccc43)cc2)ncn1)c1ccccc1 2 1 D-A-D $455.517$ $7.145$
3,3-NHC O=c1oc2ccccc2c(O)c1C(c1ccc2ccccc2c1)c1c(O)c2ccccc2oc1=O 0 0 Unknown $462.457$ $5.644$
3CP-DPS-DMAC CC1(C)c2ccccc2N(c2ccc(S(=O)(=O)c3ccc4c(c3)c3ccccc3n4-c3ccccc3)cc2)c2ccccc21 2 0 2D-0A $590.748$ $9.726$
3CP-DPS-PXZ O=S(=O)(c1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccc2c(c1)c1ccccc1n2-c1ccccc1 3 0 Multi-D/A $564.666$ $9.192$
3CzAnN c1ccc(-n2c3ccccc3c3cc(-c4c5ccccc5c(-c5ccc6ccccc6c5)c5ccccc45)ccc32)cc1 1 0 1D-0A $545.685$ $11.577$
DPQ-DMAC CC1(C)c2ccccc2N(c2ccc3nc(-c4ccccn4)c(-c4ccccn4)nc3c2)c2ccccc21 1 1 D-A $491.598$ $7.863$
DPAA c1ccc(-c2c3ccccc3c(-c3ccccc3)c3ccccc23)cc1 0 0 Unknown $330.430$ $7.327$
t3Cz-SO CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc2c(c1)c1cc(-n3c4ccc(C(C)(C)C)cc4c4cc(C(C)(C)C)ccc43)ccc1n2-c1ccc(S(=O)(=O)c2ccccc2)cc1 3 0 Multi-D/A $938.294$ $17.001$
Cz-SO O=S(=O)(c1ccccc1)c1ccc(-n2c3ccccc3c3ccccc32)cc1 1 0 1D-0A $383.472$ $5.617$
DPQ-PXZ c1ccc(-c2nc3ccc(N4c5ccccc5Oc5ccccc54)cc3nc2-c2ccccn2)nc1 2 1 D-A-D $465.516$ $7.329$
pCzAnN c1ccc2cc(-c3c4ccccc4c(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)c4ccccc34)ccc2c1 1 0 1D-0A $545.685$ $11.577$
3-BHH O=c1oc2ccccc2c(O)c1C(c1ccc(Br)s1)c1c(O)c2ccccc2oc1=O 0 0 Unknown $497.322$ $5.315$
DPP-PXZ c1ccc(-c2nc3cc(N4c5ccccc5Oc5ccccc54)cnc3nc2-c2ccccn2)nc1 1 0 1D-0A $466.504$ $6.724$
APPT-PXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2c3c(cccc13)-c1nc3c4cccnc4c4ncccc4c3nc1-2 2 1 D-A-D $537.582$ $9.102$
HITPE Oc1ccc(-c2ccc(C(=C(c3ccccc3)c3ccccc3)c3ccccc3)cc2)cc1-c1nc(-c2ccccc2)c(-c2ccccc2)n1-c1ccccc1 0 0 Unknown $718.900$ $13.253$
DV-3CzCN C=Cc1ccc(COc2cccc3c2c2ccccc2n3-c2cc(-n3c4ccccc4c4c(OCc5ccc(C=C)cc5)cccc43)c(C#N)c(-n3c4ccccc4c4c(OCc5ccc(C=C)cc5)cccc43)c2)cc1 3 2 Multi-D/A $995.195$ $17.516$
SF-DPSO O=S(=O)(c1ccc(-c2ccc3c(c2)C2(c4ccccc4-c4ccccc42)c2ccccc2-3)cc1)c1ccc(-c2ccc3c(c2)C2(c4ccccc4-c4ccccc42)c2ccccc2-3)cc1 0 0 Unknown $847.051$ $14.540$
BDQ-tBuCz CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c2nc(-c3ccccc3)c(-c3ccccc3)nc12 2 1 D-A-D $837.168$ $16.348$
BDQDMAC CC1(C)c2ccccc2N(c2ccc(N3c4ccccc4C(C)(C)c4ccccc43)c3nc(-c4ccccc4)c(-c4ccccc4)nc23)c2ccccc21 2 1 D-A-D $696.898$ $13.182$
BDQPCz c1ccc(-c2nc3c(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)ccc(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)c3nc2-c2ccccc2)cc1 2 1 D-A-D $764.932$ $14.492$
4-SCZ c1ccc2c(c1)-c1ccccc1C21c2ccccc2-c2c1ccc1[nH]c3ccccc3c21 1 0 1D-0A $405.500$ $7.665$
Py-TPA CC(C)(C)c1ccc(N(c2ccc(-c3cc4nc5c6ccc(-c7ccncc7)nc6c6nc(-c7ccncc7)ccc6c5nc4cc3C#N)cc2)c2ccc(C(C)(C)C)cc2)cc1 1 4 Multi-D/A $817.013$ $13.607$
DPACPhTPI CC(C)(C)c1ccc(-n2c(-c3ccc(-c4ccc(N5c6ccccc6C(c6ccccc6)(c6ccccc6)c6ccccc65)cc4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 1 0 1D-0A $834.079$ $16.129$
oPTBC N#Cc1ccc(-c2cc(C#N)cc(-c3ccc(C#N)cc3)c2N2c3ccccc3Oc3ccccc32)cc1 2 6 Multi-D/A $486.534$ $8.211$
mPTBC N#Cc1ccc(-c2cc(N3c4ccccc4Oc4ccccc43)cc(-c3ccc(C#N)cc3)c2C#N)cc1 2 6 Multi-D/A $486.534$ $8.211$
P3C CCCCC($C$)Cn1c2cc(C#N)ccc2c2ccc(C#Cc3ccc4ccc5cccc6ccc3c4c56)cc21 1 2 A-D-A $528.699$ $10.180$
DBCP c1ccc2c(c1)Oc1ccccc1N2c1cc2c3ccccc3c3ccc4ccc5c6ccccc6c1c1c5c4c3c21 2 0 2D-0A $531.614$ $11.648$
NAQ3 O=C1c2ccccc2C(=O)c2cc(N(c3ccc4c(c3)C(=O)c3ccccc3C4=O)c3ccc4c(c3)C(=O)c3ccccc3C4=O)ccc21 1 0 1D-0A $635.631$ $7.483$
mCP-BP-DMAC O=C(c1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccc2c(c1)c1ccccc1n2-c1cccc(-n2c3ccccc3c3ccccc32)c1 4 0 Multi-D/A $693.806$ $12.687$
BFCZPZ3 c1ccc2c(c1)oc1ccc3c4ccccc4n(-c4cnc(-n5c6ccccc6c6ccc7oc8ccccc8c7c65)cn4)c3c12 2 0 2D-0A $590.642$ $10.470$
BTCZPZ1 c1ccc2c(c1)sc1c2ccc2c3ccccc3n(-c3cnc(-n4c5ccccc5c5ccc6c7ccccc7sc6c54)cn3)c21 2 0 2D-0A $622.778$ $11.407$
BFCZPZ2 c1ccc2c(c1)oc1c2ccc2c1c1ccccc1n2-c1cnc(-n2c3ccccc3c3c4oc5ccccc5c4ccc32)cn1 2 0 2D-0A $590.642$ $10.470$
CZ9CZPZ c1ccc2c(c1)c1ccccc1n2-c1ccc2c(c1)c1ccccc1n2-c1cnc(-n2c3ccccc3c3cc(-n4c5ccccc5c5ccccc54)ccc32)cn1 4 0 Multi-D/A $740.870$ $12.865$
BFCZPZ1 c1ccc2c(c1)oc1c2ccc2c3ccccc3n(-c3cnc(-n4c5ccccc5c5ccc6c7ccccc7oc6c54)cn3)c21 2 0 2D-0A $590.642$ $10.470$
TPBPPI-PY c1ccc(-c2cc(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccc(-c5cccnc5)cc4)cc3)c(-c3ccccc3)c(-c3ccccc3)c2-c2ccccc2)cc1 0 0 Unknown $828.031$ $16.396$
TPBPPI-PBI c1ccc(-c2cc(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccc(-c5ccc(-c6nc7ccccc7n6-c6ccccc6)cc5)cc4)cc3)c(-c3ccccc3)c(-c3ccccc3)c2-c2ccccc2)cc1 0 0 Unknown 1 019.2641\,019.264 $20.007$
PTZ-2PTO O=S1(=O)c2ccccc2N(c2ccccc2)c2cc(N3c4ccccc4Sc4ccccc43)ccc21 4 0 Multi-D/A $504.636$ $8.237$
4F-m-$\nu$-DABNA Cc1cccc(N(c2cccc(C)c2)c2cc3c4c(c2)N(c2c(F)cccc2F)c2cc5c(cc2B4c2ccc(C)cc2N3c2cccc(C)c2)B2c3ccc(C)cc3N(c3cccc(C)c3)c3cc(N(c4cccc(C)c4)c4cccc(C)c4)cc(c32)N5c2c(F)cccc2F)c1 6 0 Multi-D/A 1 281.1321\,281.132 $19.815$
m-$\nu$-DABNA Cc1cccc(N(c2cccc(C)c2)c2cc3c4c(c2)N(c2ccccc2)c2cc5c(cc2B4c2ccc(C)cc2N3c2cccc(C)c2)B2c3ccc(C)cc3N(c3cccc(C)c3)c3cc(N(c4cccc(C)c4)c4cccc(C)c4)cc(c32)N5c2ccccc2)c1 6 0 Multi-D/A 1 209.1721\,209.172 $19.259$
TPA-PyI Brc1cccc(-n2c(-c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cc3)nc(-c3ccccc3)c2-c2ccccc2)c1 1 0 1D-0A $694.676$ $12.773$
4F-$\nu$-DABNA Fc1cccc(F)c1N1c2cc3c(cc2B2c4ccccc4N(c4ccccc4)c4cc(N(c5ccccc5)c5ccccc5)cc1c42)B1c2ccccc2N(c2ccccc2)c2cc(N(c4ccccc4)c4ccccc4)cc(c21)N3c1c(F)cccc1F 6 0 Multi-D/A 1 168.9161\,168.916 $17.348$
CBPTZ c1ccc(-c2nc(-c3ccccc3)nc(-c3cccc(-c4cc(-n5c6ccccc6c6ccccc65)ccc4-c4ccc(-n5c6ccccc6c6ccccc65)cc4)c3)n2)cc1 2 1 D-A-D $791.958$ $14.401$
CBPBM c1ccc(-n2c(-c3ccc(-c4cc(-n5c6ccccc6c6ccccc65)ccc4-c4ccc(-n5c6ccccc6c6ccccc65)cc4)cc3)nc3ccccc32)cc1 2 0 2D-0A $752.921$ $14.221$
CBPPY c1cncc(-c2cc(-n3c4ccccc4c4ccccc43)ccc2-c2ccc(-n3c4ccccc4c4ccccc43)cc2)c1 2 0 2D-0A $561.688$ $10.610$
N-B Cc1cc(C)c(B(c2ccccc2)c2c(C)cc(C)cc2C)c(C)c1 0 0 Unknown $326.292$ $4.053$
DPACpBr Cc1cc(N2c3ccccc3C(c3ccccc3)(c3ccccc3)c3ccccc32)cc(C)c1Br 1 0 1D-0A $516.482$ $9.232$
DPCN CC1(C)c2ccccc2N(c2ccc(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccc(C#N)cc4)cc3)cc2)c2ccccc21 1 2 A-D-A $678.839$ $12.647$
BZC-PXZ O=C(CC(=O)c1ccccc1O)c1ccc(N2c3ccccc3Oc3ccccc32)cc1 2 0 2D-0A $421.452$ $6.423$
CzBN N#Cc1c(-n2c3ccccc3c3ccccc32)cccc1-n1c2ccccc2c2ccccc21 2 2 2D-2A $433.514$ $7.752$
2CzBN CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cccc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c1C#N 2 2 2D-2A $657.946$ $12.942$
TPA-DPPI c1ccc(-c2nc3c4ccccc4c4ccccc4c3n2-c2ccc(-n3c(-c4ccc(-c5ccc(N(c6ccccc6)c6ccccc6)cc5)cc4)nc4c5ccccc5c5ccccc5c43)cc2)cc1 1 0 1D-0A $906.105$ $17.448$
ANQDC-PSTA N#Cc1cc2nc3c(nc2cc1C#N)-c1ccc(N2c4ccc(-c5ccccc5)cc4C4(c5ccccc5-c5ccccc54)c4ccc5c(sc6ccccc65)c42)c2cccc-3c12 1 5 Multi-D/A $815.961$ $14.355$
pDPAPQ-PXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2c3c(cccc13)-c1nc3cc(-c4ccncc4)c(-c4ccncc4)cc3nc1-2 2 1 D-A-D $589.658$ $10.130$
m-CZ-DPS-DMAC CC1(C)c2ccccc2N(c2cccc(S(=O)(=O)c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)c2)c2ccccc21 3 0 Multi-D/A $755.943$ $12.823$
pDPBPZ-PXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2c(c1)c1ccccc1c1nc3cc(-c4ccncc4)c(-c4ccncc4)cc3nc21 2 2 2D-2A $615.696$ $10.789$
DDPhCz-DPPZ c1ccc(-c2ccc3c(c2)c2cc(-c4ccccc4)ccc2n3-c2cc3nc4c5cccnc5c5ncccc5c4nc3cc2-n2c3ccc(-c4ccccc4)cc3c3cc(-c4ccccc4)ccc32)cc1 2 2 2D-2A $917.088$ $16.742$
DCz-DPPZ c1cnc2c(c1)c1nc3cc(-n4c5ccccc5c5ccccc54)c(-n4c5ccccc5c5ccccc54)cc3nc1c1cccnc12 2 2 2D-2A $612.696$ $10.074$
Oxd-bCZ c1ccc2c(c1)c1ccccc1n2-c1ccc(-c2ccc(-c3nnc(-c4ccc(-c5ccc(-n6c7ccccc7c7ccccc76)cc5)cc4)o3)cc2)cc1 2 0 2D-0A $704.833$ $12.932$
Fene c1ccc2c(c1)Oc1ccccc1N2c1ccc2cccc(N3c4ccccc4Oc4ccccc43)c2n1 3 0 Multi-D/A $491.550$ $9.386$
Fens c1ccc2c(c1)Sc1ccccc1N2c1ccc2cccc(N3c4ccccc4Sc4ccccc43)c2n1 3 0 Multi-D/A $523.686$ $10.104$
Yad CC1(C)c2ccccc2N(c2ccc3cccc(N4c5ccccc5C(C)(C)c5ccccc54)c3n2)c2ccccc21 1 0 1D-0A $543.714$ $10.453$
DTPCZPYTZ CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-c2ccc3ccc4c(-c5ccc(-c6nnc(-c7ccccc7)n6-c6ccccc6)cc5)ccc5ccc2c3c54)cc1 1 0 1D-0A $851.110$ $16.525$
TPADSO2 O=S1(=O)c2ccccc2S(=O)(=O)c2cc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)ccc21 1 0 1D-0A $523.635$ $6.802$
B4PyPPM c1ccc(-c2nc(-c3cc(-c4ccncc4)cc(-c4ccncc4)c3)cc(-c3cc(-c4ccncc4)cc(-c4ccncc4)c3)n2)cc1 0 1 0D-1A $616.728$ $9.726$
mAP CC1(C)c2ccccc2N(c2cccc(N3c4ccccc4C(C)(C)c4ccccc43)c2)c2ccccc21 2 0 2D-0A $492.666$ $9.905$
DBBPZ-DPXZ c1ccc(-c2cc3nc4c5ccc(N6c7ccccc7Oc7ccccc76)cc5c5cc(N6c7ccccc7Oc7ccccc76)ccc5c4nc3cc2-c2ccccc2)cc1 4 2 Multi-D/A $794.914$ $15.574$
DPIBPZ-DPXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2c(c1)c1cc(N3c4ccccc4Oc4ccccc43)ccc1c1nc3cc(-c4cncnc4)c(-c4cncnc4)cc3nc21 4 4 Multi-D/A $798.866$ $13.154$
m-DTPACO O=C(c1ccc(N(c2ccccc2)c2ccccc2)cc1)c1cccc(C(=O)c2ccc(N(c3ccccc3)c3ccccc3)cc2)c1 2 0 2D-0A $620.752$ $11.088$
mDPBPZ-DPXZ c1cncc(-c2cc3nc4c5ccc(N6c7ccccc7Oc7ccccc76)cc5c5cc(N6c7ccccc7Oc7ccccc76)ccc5c4nc3cc2-c2cccnc2)c1 4 2 Multi-D/A $796.890$ $14.364$
p-DTPACO O=C(c1ccc(C(=O)c2ccc(N(c3ccccc3)c3ccccc3)cc2)cc1)c1ccc(N(c2ccccc2)c2ccccc2)cc1 2 0 2D-0A $620.752$ $11.088$
2TDBA-SBA CC(C)(C)c1ccc2c(c1)B1c3cc(C(C)(C)C)ccc3Oc3cc(N4c5ccccc5C5(c6ccccc64)c4ccccc4N(c4cc6c7c(c4)Oc4ccc(C(C)(C)C)cc4B7c4cc(C(C)(C)C)ccc4O6)c4ccccc45)cc(c31)O2 2 0 2D-0A 1 107.0251\,107.025 $16.281$
TPA-APm c1ccc(N(c2ccccc2)c2ccc(-c3cnc4nc5c(nc4n3)-c3cccc4cccc-5c34)cc2)cc1 1 0 1D-0A $499.577$ $8.357$
TPA-AP c1ccc(N(c2ccccc2)c2ccc(-c3ccc4nc5c(nc4c3)-c3cccc4cccc-5c34)cc2)cc1 1 1 D-A $497.601$ $9.567$
TDBA-SBA CC(C)(C)c1ccc(N2c3ccccc3C3(c4ccccc42)c2ccccc2N(c2cc4c5c(c2)Oc2ccc(C(C)(C)C)cc2B5c2cc(C(C)(C)C)ccc2O4)c2ccccc23)cc1 2 0 2D-0A $858.935$ $14.256$
TPA-APy c1ccc(N(c2ccccc2)c2ccc(-c3cnc4nc5c(nc4c3)-c3cccc4cccc-5c34)cc2)cc1 1 0 1D-0A $498.589$ $8.962$
TPD4PA c1ccc(N(c2ccccc2)c2cc3c4c(c2)N(c2ccccc2)c2cc5c6c(c2B4c2ccccc2N3c2ccccc2)Oc2ccccc2B6c2ccccc2O5)cc1 3 0 Multi-D/A $779.517$ $9.967$
o-Ac2BBP CC1(C)c2ccccc2N(c2ccc(C(=O)c3ccccc3C(=O)c3ccc(N4c5ccccc5C(C)(C)c5ccccc54)cc3)cc2)c2ccccc21 2 0 2D-0A $700.882$ $12.367$
TPA-PyF2 Cc1ccc(N(c2ccc(C3(c4cccnc4)c4cc(C(C)(C)C)ccc4-c4ccc(C(C)(C)C)cc43)cc2)c2ccc(C3(c4cccnc4)c4cc(C(C)(C)C)ccc4-c4ccc(C(C)(C)C)cc43)cc2)cc1 1 0 1D-0A $966.370$ $18.171$
o-Px2BBP O=C(c1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccccc1C(=O)c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 0 Multi-D/A $648.718$ $11.300$
D-BP-PXZ O=C(c1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccccc1-c1ccccc1C(=O)c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 0 Multi-D/A $724.816$ $12.967$
BTD-PXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2nsnc2c1 2 0 2D-0A $317.373$ $5.267$
tBuCzDBA Cc1cc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)cc(C)c1B1c2ccccc2B(c2c(C)cc(-n3c4ccc(C(C)(C)C)cc4c4cc(C(C)(C)C)ccc43)cc2C)c2ccccc21 2 0 2D-0A $938.962$ $13.650$
5-Cz-Sac CC(C)(C)N1C(=O)c2cc(-n3c4ccccc4c4ccccc43)ccc2S1(=O)=O 1 0 1D-0A $404.491$ $4.727$
6-Cz-Sac CC(C)(C)N1C(=O)c2ccc(-n3c4ccccc4c4ccccc43)cc2S1(=O)=O 1 0 1D-0A $404.491$ $4.727$
DiCz-Sac CC(C)(C)N1C(=O)c2cc(-n3c4ccccc4c4ccccc43)c(-n3c4ccccc4c4ccccc43)cc2S1(=O)=O 2 0 2D-0A $569.686$ $7.824$
2CM CCCCCCCCn1c2ccccc2c2cc(Nc3nc(Nc4ccc5c(c4)c4ccccc4n5CCCCCCCC)nc(OC)n3)ccc21 2 1 D-A-D $695.956$ $12.304$
CzoB Cc1cc(C)c(B(c2ccccc2-n2c3ccccc3c3ccccc32)c2c(C)cc(C)cc2C)c(C)c1 1 0 1D-0A $491.487$ $7.150$
CzBroB Cc1cc(C)c(B(c2ccc(Br)cc2-n2c3ccccc3c3ccccc32)c2c(C)cc(C)cc2C)c(C)c1 1 0 1D-0A $570.383$ $7.913$
PxCN N#Cc1c(-c2ccc(N3c4ccccc4Oc4ccccc43)cc2)cncc1-c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 1 Multi-D/A $618.696$ $11.438$
AcCN CC1(C)c2ccccc2N(c2ccc(-c3cncc(-c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)c3C#N)cc2)c2ccccc21 2 1 D-A-D $670.860$ $12.505$
CzCNoB Cc1cc(C)c(B(c2ccc(C#N)cc2-n2c3ccccc3c3ccccc32)c2c(C)cc(C)cc2C)c(C)c1 1 2 A-D-A $516.497$ $7.022$
CNPP-TPA N#Cc1nc2c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)nc3c3nc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)ccc3c2nc1C#N 2 3 Multi-D/A $768.884$ $12.743$
SPAC-FO O=C1c2ccc(N3c4ccccc4C4(c5ccccc5-c5ccccc54)c4ccccc43)cc2-c2cc(N3c4ccccc4C4(c5ccccc5-c5ccccc54)c4ccccc43)ccc21 2 0 2D-0A $839.010$ $15.191$
DMAC-FO CC1(C)c2ccccc2N(c2ccc3c(c2)-c2cc(N4c5ccccc5C(C)(C)c5ccccc54)ccc2C3=O)c2ccccc21 2 0 2D-0A $594.758$ $11.116$
6AcBIQ CC1(C)c2ccccc2N(c2ccc3c4c(cccc24)C(=O)N(c2ccccc2)C3=O)c2ccccc21 1 0 1D-0A $480.567$ $7.750$
PtB CC(C)(C)c1ccc(-n2c(-c3ccc4ccc5cccc6ccc3c4c56)nc(-c3ccccc3)c2-c2ccccc2)cc1 0 0 Unknown $552.721$ $11.068$
BCPBT c1ccc2c(c1)c1ccccc1n2-c1cc(-c2ccc(-c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)c3nsnc23)cc(-n2c3ccccc3c3ccccc32)c1 4 0 Multi-D/A $949.155$ $17.414$
InDM CN(C)c1ccc(-n2c(-c3ccccc3)nc(-c3ccccc3)c2-c2ccccc2)cc1 0 0 Unknown $415.540$ $6.939$
PnB CCCCc1ccc(-n2c(-c3ccc4ccc5cccc6ccc3c4c56)nc(-c3ccccc3)c2-c2ccccc2)cc1 0 0 Unknown $552.721$ $11.113$
BANE c1ccc(-c2ccc(-c3c4ccccc4c(-c4c5ccccc5c(-c5ccc(-c6ccccc6)cc5)c5ccccc45)c4ccccc34)cc2)cc1 0 0 Unknown $658.844$ $14.634$
CBM CCCCn1c2ccc(C=C(C#N)C#N)cc2c2cc(C=C(C#N)C#N)ccc21 1 4 Multi-D/A $375.435$ $5.456$
DCzBNPh-1 N#Cc1ccc(-c2cccc(-c3ccc(C#N)c(-n4c5ccccc5c5ccccc54)c3)c2)cc1-n1c2ccccc2c2ccccc21 2 4 Multi-D/A $610.720$ $10.958$
MCZ-P2-DTM COc1ccc(-c2ccc3c(c2)c2cc(-c4ccc(OC)cc4)ccc2n3-c2ccc(C(=O)c3ccc4sc5ccccc5c4c3)cn2)cc1 1 0 1D-0A $666.802$ $11.129$
DBTO-PTZ O=S1(=O)c2ccc(N3c4ccccc4Oc4ccccc43)cc2-c2cc(N3c4ccccc4Oc4ccccc43)ccc21 4 0 Multi-D/A $578.649$ $9.651$
$\alpha$-DMAC-DBP CC1(C)c2ccccc2N(c2cccc3nc4c5ccccc5c5ccccc5c4nc23)c2ccccc21 1 2 A-D-A $487.606$ $9.198$
$\alpha$-2DMAC-DBP CC1(C)c2ccccc2N(c2ccc(N3c4ccccc4C(C)(C)c4ccccc43)c3nc4c5ccccc5c5ccccc5c4nc23)c2ccccc21 2 2 2D-2A $694.882$ $13.308$
$\beta$-DI-DBP c1ccc(-n2c3ccccc3c3c2c2c4ccccc4n(-c4ccc5nc6c7ccccc7c7ccccc7c6nc5c4)c2c2c4ccccc4n(-c4ccccc4)c32)cc1 3 2 Multi-D/A $775.915$ $14.381$
DIC-TRZ c1ccc(-c2nc(-c3ccccc3)nc(-n3c4ccccc4c4ccc5c6ccccc6n(-c6ccccc6)c5c43)n2)cc1 2 1 D-A-D $563.664$ $9.400$
PO-TPA CC(C)(C)c1ccc(N(c2ccc(-c3cc4c(cc3C#N)=NC3c5ccc(P(=O)(c6ccccc6)c6ccccc6)nc5-c5nc(P(=O)(c6ccccc6)c6ccccc6)ccc5C3N=4)cc2)c2ccc(C(C)(C)C)cc2)cc1 1 2 A-D-A 1 065.2111\,065.211 $13.075$
DMAC-Cz-TTR CC1(C)c2ccccc2N(c2ccc3c(c2)c2cc(N4c5ccccc5C(C)(C)c5ccccc54)ccc2n3-c2ccc3c(c2)S(=O)(=O)c2ccccc2S3(=O)=O)c2ccccc21 3 0 Multi-D/A $860.073$ $12.981$
DCzPO Cc1ccc2c(c1)c1cc(C)ccc1n2-c1cccc(-n2c3ccc(C)cc3c3cc(C)ccc32)c1P(=O)(c1ccccc1)c1ccccc1 2 0 2D-0A $664.789$ $10.754$
IQ-oTPA O=C1c2ccccc2-c2nc3cc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)c(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cc3nc21 2 1 D-A-D $718.860$ $13.115$
2C N#CC(C#N)=Cc1ccc(-c2ccc3ccc4cccc5ccc2c3c45)cc1F 0 2 0D-2A $372.402$ $6.821$
C-H N#CC(C#N)=Cc1c(F)cc(-c2ccc3ccc4cccc5ccc2c3c45)cc1F 0 2 0D-2A $390.392$ $6.960$
PSPP O=S(=O)(c1cccnc1)c1cccc(N2c3ccccc3Oc3ccccc32)c1 2 0 2D-0A $400.459$ $5.490$
2DN CC(C)c1cccc(C(C)C)c1N1C(=O)c2cc(N3c4ccccc4C(C)(C)c4ccccc43)cc3cc(N4c5ccccc5C(C)(C)c5ccccc54)cc(c23)C1=O 2 0 2D-0A $772.005$ $14.105$
PSPBP O=S(=O)(c1cccnc1)c1cc(N2c3ccccc3Oc3ccccc32)cc(N2c3ccccc3Oc3ccccc32)c1 4 0 Multi-D/A $581.653$ $9.065$
2PN CC(C)c1cccc(C(C)C)c1N1C(=O)c2cc(N3c4ccccc4Oc4ccccc43)cc3cc(N4c5ccccc5Oc5ccccc54)cc(c23)C1=O 4 0 Multi-D/A $719.841$ $13.038$
pCbz-PYR c1ccc(-c2ccc3c(c2)c2cc(-c4ccccc4)ccc2n3-c2cc(-n3c4ccc(-c5ccccc5)cc4c4cc(-c5ccccc5)ccc43)ncn2)cc1 2 1 D-A-D $714.872$ $13.339$
PPI-An-NPIM c1ccc(-c2nc(-c3ccccc3)n(-c3ccc(-c4c5ccccc5c(-c5ccc(-c6nc7c8ccccc8c8ccccc8c7n6-c6ccccc6)cc5)c5ccccc45)cc3)c2-c2ccccc2)cc1 0 0 Unknown $917.128$ $17.826$
2PS-2FPh Fc1cc(N2c3ccccc3Sc3ccccc32)c(F)cc1N1c2ccccc2Sc2ccccc21 4 0 Multi-D/A $508.618$ $9.834$
9CzTPE C(=Cc1ccc(-n2c3ccccc3c3ccccc32)cc1)c1ccc(C(=C(c2ccccc2)c2ccccc2)c2ccccc2)cc1 1 0 1D-0A $599.777$ $11.961$
2CzTPE CCn1c2ccccc2c2ccc(-c3ccc(C=Cc4ccc(C(=C(c5ccccc5)c5ccccc5)c5ccccc5)cc4)cc3)cc21 1 0 1D-0A $627.831$ $12.659$
FTAT-FBO O=P(c1ccccc1)(c1ccccc1)c1ccc([Si](c2ccccc2)(c2ccccc2)c2ccccc2)cc1 0 0 Unknown $536.687$ $4.703$
DmPy-PPI c1ccc(-n2c(-c3ccc(-c4cc(-c5cccc(-c6cccnc6)c5)cc(-c5cccc(-c6cccnc6)c5)c4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 0 0 Unknown $752.921$ $14.124$
mCP-BmPy CC(C)(C)c1ccc2c(c1)C(c1cc(-c3cccnc3)cc(-c3cccnc3)c1)(c1ccc3c(c1)c1ccccc1n3-c1cccc(-n3c4ccccc4c4ccccc43)c1)c1cc(C(C)(C)C)ccc1-2 2 0 2D-0A $915.197$ $16.963$
mCP-Py CC(C)(C)c1ccc2c(c1)C(c1cccnc1)(c1ccc3c(c1)c1ccccc1n3-c1cccc(-n3c4ccccc4c4ccccc43)c1)c1cc(C(C)(C)C)ccc1-2 2 0 2D-0A $762.013$ $14.234$
mCP-Ph CC(C)(C)c1ccc2c(c1)C(c1ccccc1)(c1ccc3c(c1)c1ccccc1n3-c1cccc(-n3c4ccccc4c4ccccc43)c1)c1cc(C(C)(C)C)ccc1-2 2 0 2D-0A $761.025$ $14.839$
DMTAC Cc1ccc(-c2c(-c3ccccc3)c(-c3ccccc3)c3c(N)cc4c5ccccc5c(N)cc4c3c2-c2ccc(C)c(C)c2)cc1C 0 0 Unknown $618.824$ $12.212$
IPTAC Cc1ccc(-c2c(-c3ccc(C(C)C)cc3)c(-c3ccc(C(C)C)cc3)c3c(c(N)cc4c5ccccc5c(N)cc43)c2-c2ccc(C)cc2)cc1 0 0 Unknown $674.932$ $13.842$
TBTAC Cc1ccccc1-c1c(-c2ccc(C(C)(C)C)cc2)c(-c2ccc(C(C)(C)C)cc2)c2c(c(N)cc3c4ccccc4c(N)cc32)c1-c1ccccc1C 0 0 Unknown $702.986$ $14.190$
p2Cz2Trz c1ccc(-c2nc(-c3ccccc3)nc(-c3cc(-n4c5ccccc5c5ccccc54)c(-c4nc(-c5ccccc5)nc(-c5ccccc5)n4)cc3-n3c4ccccc4c4ccccc43)n2)cc1 2 2 2D-2A $871.020$ $14.253$
m2Cz2Trz c1ccc(-c2nc(-c3ccccc3)nc(-c3cc(-c4nc(-c5ccccc5)nc(-c5ccccc5)n4)c(-n4c5ccccc5c5ccccc54)cc3-n3c4ccccc4c4ccccc43)n2)cc1 2 2 2D-2A $871.020$ $14.253$
2,4-CzPyCl3 Clc1nc(-n2c3ccccc3c3ccccc32)c(Cl)c(-n2c3ccccc3c3ccccc32)c1Cl 2 0 2D-0A $512.827$ $9.236$
DMAC-PCN CC1(C)c2ccccc2N(c2ccc(-c3c(C#N)c(-c4ccccc4)nc(-c4ccccc4)c3C#N)cc2)c2ccccc21 1 2 A-D-A $564.692$ $9.935$
2PXZP Cn1cnc2c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)nc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)nc21 4 1 Multi-D/A $648.726$ $10.848$
2,4,6-CzPyCl2 Clc1c(-n2c3ccccc3c3ccccc32)nc(-n2c3ccccc3c3ccccc32)c(Cl)c1-n1c2ccccc2c2ccccc21 3 0 Multi-D/A $643.577$ $11.680$
1PXZP Cn1cnc2c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)ncnc21 2 1 D-A-D $391.434$ $5.606$
3,6-T2C N#Cc1ccc(-c2nc3c4ccc(-c5ccc(N(c6ccccc6)c6ccccc6)cc5)cc4c4cc(-c5ccc(N(c6ccccc6)c6ccccc6)cc5)ccc4c3n2-c2ccc(C#N)cc2)cc1 2 4 Multi-D/A $907.093$ $17.016$
3,6-TC N#Cc1ccc(-c2nc3c4ccc(-c5ccc(N(c6ccccc6)c6ccccc6)cc5)cc4c4cc(-c5ccc(N(c6ccccc6)c6ccccc6)cc5)ccc4c3n2-c2ccccc2)cc1 2 2 2D-2A $882.083$ $17.144$
Ph-CF3 c1ccc2c(c1)c1ccccc1n2-c1ccc(Nc2ccc(-n3c4ccccc4c4ccccc43)cc2)cc1 2 0 2D-0A $499.617$ $9.624$
arylene c1ccc2c(c1)c1ccccc1n2-c1ccc(Nc2ccc(-n3c4ccccc4c4ccccc43)cc2)cc1 2 0 2D-0A $499.617$ $9.624$
TS-1 O=S(=O)(c1ccccc1)c1ccc(-n2c3ccccc3c3ccccc32)cc1 1 0 1D-0A $383.472$ $5.617$
3,6-MeCzOXD Cc1ccc2c(c1)c1cc(C)ccc1n2-c1ccccc1-c1nnc(-c2ccccc2-n2c3ccc(C)cc3c3cc(C)ccc32)o1 2 0 2D-0A $608.745$ $10.831$
p-PO15NCzDPA O=P(c1ccccc1)(c1ccccc1)c1ccc(-c2c3ccccc3c(-c3ccc(-n4c5cccnc5c5cccnc54)cc3)c3ccccc23)cc1 0 0 Unknown $697.778$ $10.853$
2,7-MeCzOXD Cc1ccc2c3ccc(C)cc3n(-c3ccccc3-c3nnc(-c4ccccc4-n4c5cc(C)ccc5c5ccc(C)cc54)o3)c2c1 2 0 2D-0A $608.745$ $10.831$
BPCN-Cz2Ph N#Cc1ccccc1-c1ccc(-n2c3ccc(-c4ccccc4)cc3c3cc(-c4ccccc4)ccc32)cc1 1 2 A-D-A $496.613$ $9.656$
o-CzOXD c1ccc(-n2c3ccccc3c3ccccc32)c(-c2nnc(-c3ccccc3-n3c4ccccc4c4ccccc43)o2)c1 2 0 2D-0A $552.637$ $9.598$
BPCN-2Cz N#Cc1ccccc1-c1ccc(-n2c3ccccc3c3cc(-c4ccc5c(c4)c4ccccc4n5-c4ccccc4)ccc32)cc1 2 2 2D-2A $585.710$ $11.086$
Cbz-MI CCCCCCN1C(=O)C(c2ccc(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)cc2)=C(c2ccc(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)cc2)C1=O 2 0 2D-0A $816.017$ $14.075$
DBFDPO O=P(c1ccccc1)(c1ccccc1)c1cccc2c1oc1c(P(=O)(c3ccccc3)c3ccccc3)cccc12 0 0 Unknown $568.549$ $6.865$
XCBP Cc1cc(-n2c3ccccc3c3ccccc32)c(C)cc1-c1ccc(C(=O)c2ccccc2)cc1 1 0 1D-0A $451.569$ $8.299$
Cbz-PYR c1ccc2c(c1)c1ccccc1n2-c1cc(-n2c3ccccc3c3ccccc32)ncn1 2 1 D-A-D $410.480$ $6.671$
tCbz-mPYRs CSc1nc(-n2c3ccc(C(C)(C)C)cc3c3c(C(C)(C)C)cccc32)c(C)c(-n2c3ccc(C(C)(C)C)cc3c3c(C(C)(C)C)cccc32)n1 2 1 D-A-D $695.033$ $12.891$
PCBP O=C(c1ccccc1)c1ccc(-c2ccc(-n3c4ccccc4c4ccccc43)cc2)cc1 1 0 1D-0A $423.515$ $7.682$
MeCZ-TRZ Cc1ccc2c(c1)c1ccccc1n2-c1ccc(-c2nc(-c3ccc(-n4c5ccccc5c5cc(C)ccc54)c(C)c3)nc(-c3ccc(-n4c5ccccc5c5cc(C)ccc54)c(C)c3)n2)cc1C 3 1 Multi-D/A $889.119$ $16.014$
3ca COc1ccc(-c2cc(-c3ccc(C(F)(F)F)cc3)c(-c3nc4ccccc4[nH]3)s2)cc1 0 0 Unknown $450.485$ $7.653$
2MeCZ-TRZ Cc1ccc2c(c1)c1cc(C)ccc1n2-c1ccc(-c2nc(-c3ccc(-n4c5ccc(C)cc5c5cc(C)ccc54)c(C)c3)nc(-c3ccc(-n4c5ccc(C)cc5c5cc(C)ccc54)c(C)c3)n2)cc1C 3 1 Multi-D/A $931.200$ $16.940$
CZ-TRZ Cc1cc(-c2nc(-c3ccc(-n4c5ccccc5c5ccccc54)c(C)c3)nc(-c3ccc(-n4c5ccccc5c5ccccc54)c(C)c3)n2)ccc1-n1c2ccccc2c2ccccc21 3 1 Multi-D/A $847.038$ $15.089$
DMAC-AQ CC1(C)c2ccccc2N(c2ccc3c(c2)C(=O)c2ccccc2C3=O)c2ccccc21 1 0 1D-0A $415.492$ $6.571$
DTPTCzDP-CN CC(C)(C)c1nc(-c2ccc(-n3c4ccc(N(c5ccccc5)c5ccccc5)cc4c4cc(N(c5ccccc5)c5ccccc5)ccc43)c(C#N)c2)nc(C(C)(C)C)n1 3 3 Multi-D/A $794.019$ $14.042$
mP2MPC Cc1cc(-c2cccc(-n3c4ccccc4c4ccccc43)c2)c(C)cc1-c1cccc(-c2nc3c4ccccc4c4ccccc4c3n2-c2ccccc2)c1 1 0 1D-0A $715.900$ $14.047$
XPhos CC(C)C1=CC(C(C)C)=C(c2ccccc2)C(C(C)C)(P(C2CCCCC2)C2CCCCC2)C1 0 0 Unknown $478.745$ $10.624$
TPXZ-as-TAZ c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2nnc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)n2)cc1 6 0 Multi-D/A $852.954$ $15.599$
2Trz6CNDBT N#Cc1cccc2c1sc1ccc(-c3nc(-c4ccccc4)nc(-c4ccccc4)n3)cc12 0 3 Multi-D/A $440.531$ $7.112$
TPAPyPI c1ccc(N(c2ccccc2)c2ccc(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4cccc(-c5ccc6ccc7cccc8ccc5c6c78)c4)cc3)cc2)cc1 1 0 1D-0A $814.004$ $16.700$
TPACzPI c1ccc(N(c2ccccc2)c2ccc(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4cccc(-c5ccc6c(c5)c5ccccc5n6-c5ccccc5)c4)cc3)cc2)cc1 2 0 2D-0A $855.057$ $16.900$
NA-TNA CC(C)(C)c1ccc(N2C(=O)c3cccc4c(/C=C/c5ccc6cc(N(c7ccccc7)c7ccccc7)ccc6c5)ccc(c34)C2=O)cc1 1 0 1D-0A $648.806$ $11.731$
CzPrSBI O=C1c2ccccc2S(=O)(=O)N1CCCn1c2ccccc2c2ccccc21 1 0 1D-0A $390.464$ $4.029$
m-CzCNDMT Cc1cc(-c2cccc(-n3c4ccccc4c4cc(C#N)ccc43)c2)c(C)cc1-c1cccc(-n2c3ccccc3c3cc(C#N)ccc32)c1 2 4 Multi-D/A $638.774$ $11.575$
DTXO-TPA2 Cc1cc(N(c2ccccc2)c2cccc3ccccc23)ccc1-c1ccc(N(c2ccccc2)c2cccc3ccccc23)cc1C 2 0 2D-0A $616.808$ $13.216$
24PQ-Cz c1ccc2c(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)nc(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)nc2c1 2 1 D-A-D $612.736$ $11.158$
SFPC O=S(=O)(c1ccc(N2c3ccccc3Sc3ccccc32)cc1)c1ccc(-n2c3ccccc3c3ccccc32)cc1 3 0 Multi-D/A $580.734$ $9.551$
2CzMC COc1cc(=O)oc2cc3c(cc12)c1ccccc1n3-c1ccccc1 1 0 1D-0A $341.366$ $4.899$
PPI-PIM-2 FC(F)(F)c1ccc(-n2c(-c3ccc(-c4ccc(-n5c(-c6ccccc6)nc(-c6ccccc6)c5-c5ccccc5)cc4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 0 0 Unknown $808.907$ $14.871$
4CzMC COc1cc(=O)oc2c1ccc1c2c2ccccc2n1-c1ccccc1 1 0 1D-0A $341.366$ $4.899$
PPI-PIM-1 FC(F)(F)c1ccc(-n2c(-c3ccc(-c4ccc(-c5nc(-c6ccccc6)c(-c6ccccc6)n5-c5ccccc5)cc4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 0 0 Unknown $808.907$ $14.871$
CzTPA c1ccc(N(c2ccccc2)c2ccccc2-n2c3ccccc3c3ccccc32)cc1 2 0 2D-0A $410.520$ $8.254$
SFX-bPy c1ccc2c(c1)Oc1ccc(-c3ccc4ccccc4n3)cc1C21c2ccccc2-c2ccccc21 0 0 Unknown $459.548$ $8.371$
t–BuCz2BP CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(C(=O)c2ccc(-n3c4ccc(C(C)(C)C)cc4c4cc(C(C)(C)C)ccc43)cc2)cc1 2 0 2D-0A $737.044$ $14.302$
SAF-2CzB CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-c2ccc3c(c2)C2(c4cc(-c5ccc(-n6c7ccc(C(C)(C)C)cc7c7cc(C(C)(C)C)ccc76)cc5)ccc4-3)c3ccccc3N(c3ccccc3)c3ccccc32)cc1 3 0 Multi-D/A 1 114.5341\,114.534 $22.551$
F-2CzB CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-c2ccc3c(c2)C(C)(C)c2cc(-c4ccc(-n5c6ccc(C(C)(C)C)cc6c6cc(C(C)(C)C)ccc65)cc4)ccc2-3)cc1 2 0 2D-0A $901.295$ $18.711$
SAF-CzB CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-c2ccc3c(c2)C2(c4ccccc4-3)c3ccccc3N(c3ccccc3)c3ccccc32)cc1 2 0 2D-0A $761.025$ $15.192$
BPy-TP2 c1ccc(-c2ccc(-c3ccc4c5ccc(-c6ccc(-c7ccccn7)nc6)cc5c5ccccc5c4c3)cn2)nc1 0 0 Unknown $536.638$ $9.394$
Cz–AQ O=C1c2ccccc2C(=O)c2cc(-n3c4ccccc4c4ccccc43)ccc21 1 0 1D-0A $373.411$ $5.559$
-ph-pCzAnBzt Cc1cc(-c2c3ccccc3c(-c3ccc(C#N)cc3)c3ccccc23)c(C)cc1-c1ccc(-n2c3ccccc3c3ccccc32)cc1 1 2 A-D-A $624.787$ $12.580$
PyI c1cc2ccc3cccc4c5[nH]cnc5c(c1)c2c34 0 0 Unknown $242.281$ $4.460$
DVCz-2CzCN C=Cc1ccc(COc2cccc3c2c2ccccc2n3CCCCCCOc2cccc3c2c2ccccc2n3-c2cccc(-n3c4ccccc4c4c(OCCCCCCn5c6ccccc6c6c(OCc7ccc(C=C)cc7)cccc65)cccc43)c2C#N)cc1 4 2 Multi-D/A 1 228.5501\,228.550 $21.691$
ATDBT c1ccc(-c2c3ccccc3c(-c3ccc(-n4c5ccccc5c5c6sc7ccccc7c6ccc54)cc3)c3ccccc23)cc1 1 0 1D-0A $601.774$ $12.792$
DPhBCz c1ccc(-n2c3ccccc3c3cc(-c4ccc5c(c4)c4ccccc4n5-c4ccccc4)ccc32)cc1 2 0 2D-0A $484.602$ $9.548$
AcHPM CC1(C)c2ccccc2N(c2ccc(-c3cc(-c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)ncn3)cc2)c2ccccc21 2 1 D-A-D $646.838$ $12.029$
PhCO-DMAC CC1(C)c2ccccc2N(c2ccc(C(=O)c3ccccc3)cc2)c2ccccc21 1 0 1D-0A $389.498$ $7.027$
Diphenylsulfane c1ccc(-c2cccc(-c3nc(-c4cccc(-c5ccccc5)c4)nc(-c4cccc(-c5ccccc5)c4)n3)c2)cc1 0 1 0D-1A $537.666$ $9.874$
2BrCPT Brc1ccc2c(c1)c1cc(Br)ccc1n2-c1ccc(-c2nc(-c3ccccc3)nc(-c3ccccc3)n2)cc1 1 1 D-A $632.359$ $9.495$
DCbz-Ph-Bz c1ccc(-n2c(-c3ccc(-c4cc(-n5c6ccccc6c6ccccc65)cc(-n5c6ccccc6c6ccccc65)c4)cc3)nc3ccccc32)cc1 2 0 2D-0A $676.823$ $12.554$
DDPA-Ph-Bz c1ccc(N(c2ccccc2)c2cc(-c3ccc(-c4nc5ccccc5n4-c4ccccc4)cc3)cc(N(c3ccccc3)c3ccccc3)c2)cc1 2 0 2D-0A $680.855$ $13.299$
IDAC-MCO Cc1cc2cc(-n3c4ccccc4c4cc5c(cc43)C(C)(C)c3ccccc3N5c3ccccc3)ccc2c(=O)o1 2 0 2D-0A $532.643$ $9.308$
diarylbiindole c1ccc(-c2c(-c3ccccc3)c3ccc4cccc5cc(-c6ccccc6)c(c2-c2ccccc2)c3c45)cc1 0 0 Unknown $506.648$ $11.252$
CNDSB N#CC(=Cc1ccccc1)c1ccc(C(C#N)=Cc2ccccc2)cc1 0 2 0D-2A $332.406$ $5.815$
TMAB Cc1cc(C)c(CC#Cc2ccc(N(C)C)cc2)c(B(c2c(C)c(C)cc(C)c2CC#Cc2ccc(N(C)C)cc2)c2c(C)c(C)cc(C)c2CC#Cc2ccc(N(C)C)cc2)c1C 0 0 Unknown $840.021$ $9.959$
MOF c1cncc(CC(c2nc3ccccc3[nH]2)N(C(Cc2cccnc2)c2nc3ccccc3[nH]2)C(Cc2cccnc2)c2nc3ccccc3[nH]2)c1 0 0 Unknown $680.820$ $8.054$
DICzTrz N#Cc1cc(-c2nc(-c3ccccc3)nc(-c3ccccc3)n2)ccc1-n1c2ccccc2c2cc3c4ccccc4n4c5ccccc5c(c21)c34 3 3 Multi-D/A $662.756$ $10.991$
PXZ2PTO O=S1(=O)c2ccccc2N(c2ccccc2)c2cc(N3c4ccccc4Oc4ccccc43)ccc21 4 0 Multi-D/A $488.568$ $7.878$
SPBP-DPAC O=C1c2ccccc2C2(c3cc(N4c5ccccc5C(c5ccccc5)(c5ccccc5)c5ccccc54)ccc31)c1ccccc1-c1ccccc12 1 0 1D-0A $675.831$ $11.760$
PX-TRZ Cc1cc(C)c(B(c2ccc(-c3nc(Cl)nc(-c4ccccc4)n3)cc2)c2c(C)cc(C)cc2C)c(C)c1 0 1 0D-1A $515.897$ $6.226$
C2MPI Cc1ccc(N(c2ccc(C)cc2)c2ccccc2C2(c3ccccc3N(c3ccc(C)cc3)c3ccc(C)cc3)CCCCC2)cc1 2 0 2D-0A $626.888$ $13.110$
p-PCzTPD CCCCCCCCN1C(=O)c2c(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)sc(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)c2C1=O 2 0 2D-0A $747.964$ $13.233$
m-PCzTPD CCCCCCCCN1C(=O)c2c(-c3cccc(-n4c5ccccc5c5ccccc54)c3)sc(-c3cccc(-n4c5ccccc5c5ccccc54)c3)c2C1=O 2 0 2D-0A $747.964$ $13.233$
o-PCzTPD CCCCCCCCN1C(=O)c2c(-c3ccccc3-n3c4ccccc4c4ccccc43)sc(-c3ccccc3-n3c4ccccc4c4ccccc43)c2C1=O 2 0 2D-0A $747.964$ $13.233$
CPBzFC c1ccc2c(c1)c1ccccc1n2-c1ccc(-c2ccc(-c3ccc4c(c3)Cc3cc(-n5c6ccccc6c6ccccc65)ccc3-4)c3nnsc23)cc1 2 0 2D-0A $706.874$ $12.791$
SiCz2Py2 c1cncc(-c2ccc([Si](c3ccc(-c4cccnc4)cc3)(c3ccc(-n4c5ccccc5c5ccccc54)cc3)c3ccc(-n4c5ccccc5c5ccccc54)cc3)cc2)c1 2 0 2D-0A $821.072$ $11.382$
SiCz1Py3 c1cncc(-c2ccc([Si](c3ccc(-c4cccnc4)cc3)(c3ccc(-c4cccnc4)cc3)c3ccc(-n4c5ccccc5c5ccccc54)cc3)cc2)c1 1 0 1D-0A $732.963$ $9.347$
SiCz3Py1 c1cncc(-c2ccc([Si](c3ccc(-n4c5ccccc5c5ccccc54)cc3)(c3ccc(-n4c5ccccc5c5ccccc54)cc3)c3ccc(-n4c5ccccc5c5ccccc54)cc3)cc2)c1 3 0 Multi-D/A $909.181$ $13.417$
POBBPE CC(C)(C)c1ccc(Oc2ccc(C(C)(C)C)cc2P(=O)(c2ccccc2)c2ccccc2)cc1 0 0 Unknown $482.604$ $7.713$
CzppT c1ccc(-c2c(-c3ccccc3)c(-c3ccc(-c4cccc5c4[nH]c4ccccc45)cc3)c(-c3ccccn3)c(-c3ccccn3)c2-c2ccc(-c3cccc4c3[nH]c3ccccc34)cc2)cc1 2 0 2D-0A $867.068$ $17.082$
SBDBQ-PXZ c1ccc(-c2nc3ccc(N4c5ccccc5Oc5ccccc54)cc3nc2-c2ccccc2)cc1 2 1 D-A-D $463.540$ $8.539$
DBQ-3PXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2nc3ccc(N4c5ccccc5Oc5ccccc54)cc3nc2-c2ccc(N3c4ccccc4Oc4ccccc43)cc2)cc1 6 1 Multi-D/A $825.928$ $15.690$
P4CzCN–BCz C=Cc1ccc(COc2ccc(C(C)(C)c3ccc(O)cc3)cc2)cc1 0 0 Unknown $344.454$ $5.940$
NAP-1,5-DPA COc1ccc(N(c2ccc(OC)cc2)c2cccc3c(N(c4ccc(OC)cc4)c4ccc(OC)cc4)cccc23)cc1 2 0 2D-0A $582.700$ $9.814$
3DPA3CN N#Cc1c(-c2ccc(N(c3ccccc3)c3ccccc3)cc2)c(C#N)c(-c2ccc(N(c3ccccc3)c3ccccc3)cc2)c(C#N)c1-c1ccc(N(c2ccccc2)c2ccccc2)cc1 3 6 Multi-D/A $883.071$ $16.712$
TPA-AN-NA c1ccc(N(c2ccccc2)c2ccc(-c3c4ccccc4c(-c4ccc5ccccc5c4)c4ccccc34)cc2)cc1 1 0 1D-0A $547.701$ $11.950$
OCzBN COc1ccc2c(c1)c1ccccc1n2-c1cc(-n2c3ccccc3c3cc(OC)ccc32)c(C#N)c(-n2c3ccccc3c3cc(OC)ccc32)c1 3 2 Multi-D/A $688.787$ $10.875$
BBCz c1ccc2c(c1)[nH]c1cc3[nH]c4ccccc4c3cc12 2 0 2D-0A $256.308$ $4.956$
IQ-Se COc1c(C#N)c(C)cc2c([Se][Se]c3c(C)nc(N4CCCCC4)c4c(OC)c(C#N)c(C)cc34)c(C)nc(N3CCCCC3)c12 0 4 Multi-D/A $746.676$ $5.032$
dTolmp Cc1ccccc1P(CP(CP(c1ccccc1C)c1ccccc1C)c1ccccc1)c1ccccc1C 0 0 Unknown $562.614$ $8.208$
SiTCNCz N#Cc1cccc(-c2nc(-c3cccc([Si](c4ccccc4)(c4ccccc4)c4ccccc4)c3)nc(-n3c4ccccc4c4ccccc43)n2)c1 1 3 Multi-D/A $681.875$ $7.552$
dpmb Cc1cc(P(c2ccccc2)c2ccccc2)c(P(c2ccccc2)c2ccccc2)cc1C 0 0 Unknown $474.524$ $5.820$
3CzBN N#Cc1c(-n2c3ccccc3c3ccccc32)cc(-n2c3ccccc3c3ccccc32)cc1-n1c2ccccc2c2ccccc21 3 2 Multi-D/A $598.709$ $10.850$
TCB O=C(c1ccccc1)c1ccc(-n2c3ccccc3c3cc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)ccc32)cc1 2 0 2D-0A $590.726$ $11.151$
CPzPyC c1ccc2c(c1)c1ccccc1n2-c1cncc(-c2cncc(-n3c4ccccc4c4ccccc43)n2)c1 2 0 2D-0A $487.566$ $7.733$
CPzPzC c1ccc2c(c1)c1ccccc1n2-c1cncc(-c2cncc(-n3c4ccccc4c4ccccc43)n2)n1 2 0 2D-0A $488.554$ $7.128$
CPzPC c1cc(-c2cncc(-n3c4ccccc4c4ccccc43)n2)cc(-n2c3ccccc3c3ccccc32)c1 2 0 2D-0A $486.578$ $8.338$
CBIQD O=C1c2cccc3cccc(c23)C(=O)N1c1nc2ccc(Cl)cc2s1 0 0 Unknown $364.813$ $4.904$
PXZ-QL c1ccc2c(c1)Oc1ccccc1N2c1ccc2ccccc2n1 1 0 1D-0A $310.356$ $5.810$
SBA c1ccc2c(c1)Nc1ccccc1C21c2ccccc2Nc2ccccc21 0 0 Unknown $346.433$ $6.183$
TATP-BP O=C(c1ccc(-n2c3ccccc3c3c4c(c5ccccc5n4-c4ccccc4)c4c(c5ccccc5n4-c4ccccc4)c32)cc1)c1ccc(-n2c3ccccc3c3c4c(c5ccccc5n4-c4ccccc4)c4c(c5ccccc5n4-c4ccccc4)c32)cc1 6 0 Multi-D/A 1 173.3921\,173.392 $21.500$
Pd-1 CCCCN(CCCC)c1ccc(C=Nc2cc(C#N)c(C#N)cc2N=Cc2ccc(N(CCCC)CCCC)cc2O)c(O)c1 0 4 Multi-D/A $620.842$ $9.156$
C3F7 Cc1cc(C)c(B(c2ccccc2-n2c3ccccc3c3ccccc32)c2c(C)cc(C)cc2C)c(C)c1 1 0 1D-0A $491.487$ $7.150$
BuCzCF3oB Cc1cc(C)c(B(c2ccc(C(F)(F)F)cc2-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c2c(C)cc(C)cc2C)c(C)c1 1 0 1D-0A $671.700$ $10.764$
phenazasiline Brc1ccc2nc3c4ccc(Br)cc4c4ccccc4c3nc2c1 0 2 0D-2A $438.122$ $6.614$
D–A–D Brc1ccc2nc3c4ccc(Br)cc4c4ccccc4c3nc2c1 0 2 0D-2A $438.122$ $6.614$
pzpy c1ccc(-c2nc3ccccc3n2-c2ccc(-c3c4ccccc4c(-c4ccc(-n5c(-c6ccccc6)nc6ccccc65)cc4)c4ccccc34)cc2)cc1 0 0 Unknown $714.872$ $13.339$
m-PO-ABN N#Cc1ccc(-c2c3ccccc3c(-c3cccc(-c4nc5c6ccccc6c6ccccc6c5o4)c3)c3ccccc23)cc1 0 2 0D-2A $572.667$ $11.313$
PO-ABN N#Cc1ccc(-c2c3ccccc3c(-c3cccc(-c4nc5c6ccccc6c6ccccc6c5o4)c3)c3ccccc23)cc1 0 2 0D-2A $572.667$ $11.313$
p-PO-ABN N#Cc1ccc(-c2c3ccccc3c(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5o4)cc3)c3ccccc23)cc1 0 2 0D-2A $572.667$ $11.313$
teracridan c1ccc2c(c1)Cc1c(cccc1-c1ccc3c(c1-c1cccc4c1Cc1ccccc1N4)Cc1ccccc1N3)N2 0 0 Unknown $539.682$ $9.964$
PTPC CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-c2ccc(-c3cc(-c4ccccc4)c(-c4ccc(-c5nc6c7ccccc7c7ccccc7c6n5-c5ccccc5)cc4)cc3-c3ccccc3)cc2)cc1 1 0 1D-0A 1 028.3561\,028.356 $21.026$
TPTI O=c1[nH]c2ccsc2c2cc3c(=O)[nH]c4ccsc4c3cc12 0 0 Unknown $324.386$ $3.799$
BDBTCz-Trz c1ccc(-c2nc(-c3ccccc3)nc(-c3ccc(-n4c5ccc(-c6ccc7sc8ccccc8c7c6)cc5c5cc(-c6ccc7sc8ccccc8c7c6)ccc54)cc3)n2)cc1 1 1 D-A $839.061$ $16.039$
BDBFCz-Trz c1ccc(-c2nc(-c3ccccc3)nc(-c3ccc(-n4c5ccc(-c6ccc7oc8ccccc8c7c6)cc5c5cc(-c6ccc7oc8ccccc8c7c6)ccc54)cc3)n2)cc1 1 1 D-A $806.925$ $15.102$
HAP-3MF Cc1cc(C2=NC3=NC(c4ccc(F)c(C)c4)=NC4=NC(c5ccc(F)c(C)c5)=NC(=N2)N34)ccc1F 0 0 Unknown $497.484$ $5.080$
FlCz CC1(C)c2ccccc2-c2ccc(-c3ncc(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)cn3)cc21 1 1 D-A $513.644$ $9.214$
SFCz c1ccc2c(c1)-c1ccccc1C21c2ccccc2-c2ccc(-c3ncc(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)cn3)cc21 1 1 D-A $635.770$ $11.251$
o-TPA-AP-TPA c1ccc(N(c2ccccc2)c2ccccc2-c2c3ccccc3c(-c3ccc4ccc5c(-c6ccccc6N(c6ccccc6)c6ccccc6)ccc6ccc3c4c65)c3ccccc23)cc1 2 0 2D-0A $865.092$ $18.831$
p-TPA-AP-TPA c1ccc(N(c2ccccc2)c2ccc(-c3c4ccccc4c(-c4ccc5ccc6c(-c7ccc(N(c8ccccc8)c8ccccc8)cc7)ccc7ccc4c5c76)c4ccccc34)cc2)cc1 2 0 2D-0A $865.092$ $18.831$
TNPZ N#Cc1ccc(-c2ccc3c(c2)c2cc(-c4ccc(C#N)cc4)ccc2c2nc4cc(-c5ccc(N(c6ccccc6)c6ccccc6)cc5)c(-c5ccc(N(c6ccccc6)c6ccccc6)cc5)cc4nc32)cc1 2 6 Multi-D/A $969.164$ $18.440$
m-TPA-AP-TPA c1ccc(N(c2ccccc2)c2cccc(-c3c4ccccc4c(-c4ccc5ccc6c(-c7cccc(N(c8ccccc8)c8ccccc8)c7)ccc7ccc4c5c76)c4ccccc34)c2)cc1 2 0 2D-0A $865.092$ $18.831$
5,6PXZ-PIDO CC1(C)C(=O)c2cc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)cc2C1=O 4 0 Multi-D/A $688.783$ $12.577$
TRZ-3SO2 O=S(=O)(c1ccccc1)c1cccc(-c2nc(-c3cccc(S(=O)(=O)c4ccccc4)c3)nc(-c3cccc(S(=O)(=O)c4ccccc4)c3)n2)c1 0 1 0D-1A $729.861$ $7.371$
35tCzBPym CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cc(C(=O)c2cncnc2)cc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c1 2 1 D-A-D $739.020$ $13.092$
35CzBPym O=C(c1cncnc1)c1cc(-n2c3ccccc3c3ccccc32)cc(-n2c3ccccc3c3ccccc32)c1 2 1 D-A-D $514.588$ $7.902$
25CzBPym O=C(c1cncnc1)c1cc(-n2c3ccccc3c3ccccc32)ccc1-n1c2ccccc2c2ccccc21 2 1 D-A-D $514.588$ $7.902$
25tCzBPym CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c(C(=O)c2cncnc2)c1 2 1 D-A-D $739.020$ $13.092$
BPBiPA c1ccc(-c2nc3ccccc3n2-c2ccc(-c3c4ccccc4c(-c4ccc(-n5c(-c6ccccc6)nc6ccccc65)cc4)c4ccccc34)cc2)cc1 0 0 Unknown $714.872$ $13.339$
BTZ–DMAC CC1(C)c2ccccc2N(c2ccc(N3c4ccccc4C(C)(C)c4ccccc43)c3nsnc23)c2ccccc21 2 0 2D-0A $550.731$ $9.909$
BTZ–DPA c1ccc(N(c2ccccc2)c2ccc(N(c3ccccc3)c3ccccc3)c3nsnc23)cc1 2 0 2D-0A $470.601$ $8.631$
2F c1ccc(N(c2ccccc2)c2ccc3c(c2)c2cc(N(c4ccccc4)c4ccccc4)ccc2n3-c2ccc(-c3ccc4nccnc4c3)cc2)cc1 3 1 Multi-D/A $705.865$ $13.333$
PFBP-2a CC(C)(C)c1ccc2c(c1)C(C)(C)c1cc(C(C)(C)C)ccc1N2c1c($F)c($-c2c(F)c(F)c(F)c(F)c2F)c(F)c1F 1 0 1D-0A $635.614$ $11.310$
PFBP-2b CC(C)(C)c1ccc2c(c1)C(C)(C)c1cc(C(C)(C)C)ccc1N2c1c($F)c($-c2c(F)c(F)c(N3c4ccc(C(C)(C)C)cc4C(C)(C)c4cc(C(C)(C)C)ccc43)c(F)c2F)c(F)c1F 2 0 2D-0A $937.116$ $17.875$
PXZ-PPO Cc1cc2nc(N3c4ccccc4Oc4ccccc43)ccc2c(=O)o1 1 0 1D-0A $342.354$ $5.072$
PXZ-BOO Cc1nc2cc(N3c4ccccc4Oc4ccccc43)ccc2c(=O)o1 2 0 2D-0A $342.354$ $5.072$
PPZTPI CC(C)(C)c1ccc(-n2c(-c3ccc(N4c5ccccc5N(c5ccccc5)c5ccccc54)cc3)nc3c4ccccc4c4ccccc4c32)cc1 2 0 2D-0A $682.871$ $13.549$
PPZPPI c1ccc(N2c3ccccc3N(c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccccc4)cc3)c3ccccc32)cc1 2 0 2D-0A $626.763$ $12.252$
Pra-2DMAC CC1(C)c2ccccc2N(c2ccnc(N3c4ccccc4C(C)(C)c4ccccc43)c2)c2ccccc21 0 0 Unknown $493.654$ $9.300$
BAPCN CC1(C)c2ccccc2N(c2ccccc2)c2ccc3cc(-c4ccc(C#N)cc4)ccc3c21 1 2 A-D-A $436.558$ $8.488$
BACN CC1(C)c2ccccc2N(c2ccccc2)c2ccc3cc(C#N)ccc3c21 1 2 A-D-A $360.460$ $6.821$
BDQC-2 c1ccc(-c2ccc(N(c3ccc(-c4ccccc4)cc3)c3ccc4c5ccccc5n(-c5ccc6nc(-c7ccccc7)c(-c7ccccc7)nc6c5)c4c3)cc2)cc1 2 1 D-A-D $766.948$ $14.865$
NIDPA-1 c1ccc(N(c2ccccc2)c2ccc3c(c2)cc2c4cccc(N(c5ccccc5)c5ccccc5)c4n4c5ccccc5c3c24)cc1 4 0 Multi-D/A $625.775$ $12.929$
NIDPA-2 c1ccc(N(c2ccccc2)c2ccc3c(c2)c2cc(N(c4ccccc4)c4ccccc4)cc4c5c6ccccc6ccc5n3c24)cc1 4 0 Multi-D/A $625.775$ $12.929$
SABIQ c1ccc(N2c3ccccc3C3(c4ccccc4-c4nc5cc6ccccc6cc5nc43)c3ccccc32)cc1 1 1 D-A $509.612$ $8.929$
Perylene c1cc2cccc3c4cccc5cccc(c(c1)c23)c54 0 0 Unknown $252.316$ $5.737$
SAIP c1ccc(N2c3ccccc3C3(c4ccccc4-c4nccnc43)c3ccccc32)cc1 1 0 1D-0A $409.492$ $6.623$
FAC c1ccc2c(c1)Nc1ccccc1C21c2ccccc2-c2ccccc21 0 0 Unknown $331.418$ $6.107$
phenylperylene c1ccc(-c2ccc3cccc4c5cccc6cccc(c2c34)c65)cc1 0 0 Unknown $328.414$ $7.404$
ImPy-2 COc1ccc2cc(-c3nc(-c4ccccc4)c4ccccn34)ccc2c1 0 0 Unknown $350.421$ $5.830$
ImPy-1 c1ccc(-c2nc(-c3ccc4ccccc4c3)n3ccccc23)cc1 0 0 Unknown $320.395$ $5.822$
ImPy-3 c1ccc(-c2nc(-c3ccc4ccc5cccc6ccc3c4c56)n3ccccc23)cc1 0 0 Unknown $394.477$ $7.566$
BTaz-Ar-PXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2nc3ccccc3s2)cc1 2 0 2D-0A $392.483$ $7.539$
FSF4A CC1(C)c2ccccc2-c2ccc(N(c3ccccc3-c3ccccc3)c3cccc4c3-c3ccccc3C43c4ccccc4-c4ccccc43)cc21 1 0 1D-0A $675.875$ $13.473$
TPA-L-BN N#Cc1ccc(CCCc2ccc(N(c3ccccc3)c3ccccc3)cc2)cc1 1 2 A-D-A $388.514$ $7.203$
IpCm-PhBzAc CC(C)c1ccc2occ(-c3ccc(N4c5ccccc5C(C)(C)c5cc6c(cc54)oc4ccccc46)cc3)c(=O)c2c1 1 0 1D-0A $561.681$ $10.592$
DphAn-Br O=C1c2ccccc2C(c2ccccc2)(c2ccc(Br)cc2)c2ccccc21 0 0 Unknown $425.325$ $6.376$
5-DBFBI c1ccc2c(c1)nc1n(-c3ccc(-c4ccc5oc6ccccc6c5c4)cn3)c3ccccc3n21 0 0 Unknown $450.501$ $7.393$
4-DBFBI c1ccc2c(c1)nc1n(-c3cc(-c4ccc5oc6ccccc6c5c4)ccn3)c3ccccc3n21 0 0 Unknown $450.501$ $7.393$
DMACMNPTO Cc1cc(N2c3ccccc3S(=O)(=O)c3ccccc32)c(C)cc1N1c2ccccc2C(C)(C)c2ccccc21 3 0 Multi-D/A $542.704$ $9.029$
1CzSO CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc2c(c1)-c1cc(-n3c4ccc(C(C)(C)C)cc4c4cc(C(C)(C)C)ccc43)ccc1S2(=O)=O 2 0 2D-0A $771.083$ $13.884$
DpAn-BzAc CC1(C)c2ccccc2N(c2ccc(C3(c4ccccc4)c4ccccc4C(=O)c4ccccc43)cc2)c2c1ccc1c2oc2ccccc21 1 0 1D-0A $643.786$ $11.622$
BzITz N#Cc1cc(-c2nc(-c3ccccc3)nc(-c3ccccc3)n2)ccc1-n1c2ccccc2n2c3ccccc3nc12 0 3 Multi-D/A $539.602$ $7.489$
TPXZ–as-TAZ c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2nnc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)n2)cc1 6 0 Multi-D/A $852.954$ $15.599$
TPA-CN-N4-2PY CC(C)(C)c1ccc(N(c2ccc(-c3cc(N)c(N)cc3C#N)cc2)c2ccc(C(C)(C)C)cc2)cc1 1 2 A-D-A $488.679$ $8.454$
TPA-CN-N4 c1ccc2nc3c4cccnc4c4ncccc4c3nc2c1 0 2 0D-2A $282.306$ $3.879$
6,7-DCNQx-DICz N#Cc1cc2nc(-c3ccccc3)c(-c3ccc(-n4c5ccccc5c5c6c(c7ccccc7n6-c6ccccc6)c6c(c7ccccc7n6-c6ccccc6)c54)cc3)nc2cc1C#N 3 5 Multi-D/A $827.951$ $13.998$
PBICT c1ccc(-c2nc(-n3c4ccccc4c4ccc5c6ccccc6n(-c6ccccc6)c5c43)nc(-n3c4ccccc4c4ccc5c6ccccc6n(-c6ccccc6)c5c43)n2)cc1 4 1 Multi-D/A $817.956$ $13.927$
DCBNPE O=[N+]([O-])c1ccc(/C(=C(/c2ccc([N+](=O)[O-])cc2)n2c3ccccc3c3ccccc32)n2c3ccccc3c3ccccc32)cc1 2 0 2D-0A $600.634$ $9.641$
TPA-BT-Q Cc1ccc(N(c2ccc(C)cc2)c2ccc(C3(c4ccc(N(c5ccc(C)cc5)c5ccc(C)cc5)cc4)CCCCC3)cc2)cc1 2 0 2D-0A $626.888$ $13.110$
D5 c1cc(-c2ccc(-n3c4ccccc4c4ccccc43)cc2)cc(-c2ccc3c4ccccc4c4ccccc4c3c2)c1 1 0 1D-0A $545.685$ $11.577$
D6 c1cc(-c2ccc3c4ccccc4c4ccccc4c3c2)cc(-n2c3ccccc3c3ccccc32)c1 1 0 1D-0A $469.587$ $9.910$
PxPYM c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2cnc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)nc2)cc1 4 1 Multi-D/A $594.674$ $10.962$
D1–D6 c1cc(-c2ccc3c4ccccc4c4ccccc4c3c2)cc(-n2c3ccccc3c3ccccc32)c1 1 0 1D-0A $469.587$ $9.910$
DTPCZTZ CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-c2ccc(-c3nnc(-c4ccccc4)n3-c3ccccc3)cc2)cc1 1 0 1D-0A $650.870$ $11.960$
DHPZ-2BI c1ccc(-n2c(-c3ccc(N4c5ccccc5N(c5ccc(-c6nc7ccccc7n6-c6ccccc6)cc5)c5ccccc54)cc3)nc3ccccc32)cc1 2 0 2D-0A $718.864$ $12.951$
DHPZ-2BN N#Cc1ccc(N2c3ccccc3N(c3ccc(C#N)cc3)c3ccccc32)cc1 2 4 Multi-D/A $384.442$ $6.683$
SAF-2NP c1ccc2c(c1)-c1ccccc1C21c2ccccc2N(c2ccc3nc4c5cccnc5c5ncccc5c4nc3c2)c2ccccc21 1 2 A-D-A $611.708$ $10.026$
BTPO O=S1(=O)c2ccccc2N2c3ccccc3S(=O)(=O)c3cccc1c32 3 0 Multi-D/A $369.423$ $3.449$
PyDCN–DMAC CC1(C)c2ccccc2N(c2ccc(-c3cc(-c4ccc(C#N)cc4)nc(-c4ccc(C#N)cc4)c3)cc2)c2ccccc21 1 4 Multi-D/A $564.692$ $9.935$
PyDCN N#Cc1ccc(-c2cc(-c3ccccc3)cc(-c3ccc(C#N)cc3)n2)cc1 0 4 Multi-D/A $357.416$ $5.826$
PXZ-OXD c1ccc(-c2nnc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)o2)cc1 2 0 2D-0A $403.441$ $6.979$
TCz-3PA-TCz CC(C)(C)c1ccc(-c2ccc3c(c2)c2cc(-c4cnc(-c5ccc6c(c5)c5cc(-c7ccc(C(C)(C)C)cc7)ccc5n6-c5ccc(C(C)(C)C)cc5)cn4)ccc2n3-c2ccc(C(C)(C)C)cc2)cc1 2 0 2D-0A $939.304$ $18.529$
NAI–BiFA Cc1c(-c2ccccc2)cc(N2C(=O)c3cccc4c(N(c5ccc(-c6ccccc6)cc5)c5ccc6c(c5)C(C)(C)c5ccccc5-6)ccc(c34)C2=O)cc1-c1ccccc1 1 0 1D-0A $798.986$ $14.726$
NAI–PhBiFA Cc1c(-c2ccccc2)cc(N2C(=O)c3cccc4c(-c5ccc(N(c6ccc(-c7ccccc7)cc6)c6ccc7c(c6)C(C)(C)c6ccccc6-7)cc5)ccc(c34)C2=O)cc1-c1ccccc1 1 0 1D-0A $875.084$ $16.393$
pDTCz-2DPyS CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(S(=O)(=O)c2ccc(-n3c4ccc(C(C)(C)C)cc4c4cc(C(C)(C)C)ccc43)cn2)nc1 2 0 2D-0A $775.075$ $12.694$
2SPBAC-BP O=C(c1ccc(N2c3ccccc3C3(c4ccccc4-c4ccccc43)c3cc4ccccc4cc32)cc1)c1ccc(N2c3ccccc3C3(c4ccccc4-c4ccccc43)c3cc4ccccc4cc32)cc1 2 0 2D-0A $941.146$ $17.517$
SPBAC-PH c1ccc(N2c3ccccc3C3(c4ccccc4-c4ccccc43)c3cc4ccccc4cc32)cc1 1 0 1D-0A $457.576$ $8.986$
DFPBI CC1(C)c2cc(-c3nc(-c4ccccc4)c(-c4ccccc4)n3-c3ccccc3)ccc2-c2ccc(N(c3ccccc3)c3ccccc3)cc21 1 0 1D-0A $655.845$ $12.649$
TDPA–TRZ c1ccc(N(c2ccccc2)c2ccc(-c3nc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)nc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)n3)cc2)cc1 3 1 Multi-D/A $811.005$ $15.282$
DMAC-11-DPPZ CC1(C)c2ccccc2N(c2ccc3nc4c5ccccc5c5ccccc5c4nc3c2)c2ccccc21 1 2 A-D-A $487.606$ $9.198$
PXZ-11-DPPZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2nc3c4ccccc4c4ccccc4c3nc2c1 2 2 2D-2A $461.524$ $8.665$
PXZ-10-DPPZ c1ccc2c(c1)Oc1ccccc1N2c1cccc2nc3c4ccccc4c4ccccc4c3nc12 2 2 2D-2A $461.524$ $8.665$
CCC c1ccc2c(c1)Cc1c-2ccc2c1-c1ccccc1C2 0 0 Unknown $254.332$ $4.829$
CzmPPC c1ccc(-c2cccc(-c3nc(-c4cccc(-c5ccccc5)c4)nc(-c4cccc(-c5ccccc5)c4)n3)c2)cc1 0 1 0D-1A $537.666$ $9.874$
C2FLA-1 CCCC1(CCC)c2ccccc2-c2ccc(C#Cc3ccc4c(c3)C(CCC)(CCC)c3cc(C#N)ccc3-4)cc21 0 2 0D-2A $547.786$ $10.691$
MDP3FL Cc1cccc(N(c2ccccc2)c2ccc3c(c2)C(C)(C)c2cc(-c4ccc5c(c4)C(C)(C)c4cc(-c6ccc7c(c6)C(C)(C)c6cc(N(c8ccccc8)c8cccc(C)c8)ccc6-7)ccc4-5)ccc2-3)c1 2 0 2D-0A $941.275$ $19.496$
C3FLA-2 CCCCC1(CCCC)c2cc(C#Cc3ccc4c(c3)C(CCC)(CCC)c3cc(C#N)ccc3-4)ccc2-c2ccc(C#Cc3ccc4c(c3)C(CCC)(CCC)c3cc(C#N)ccc3-4)cc21 0 4 Multi-D/A $873.241$ $16.610$
9CzFDPEPO O=P(c1ccccc1)(c1ccccc1)c1ccccc1Oc1ccc(-c2ccc(C3(c4ccc(-n5c6ccccc6c6ccccc65)cc4)c4ccccc4-c4ccccc43)cc2)cc1P(=O)(c1ccccc1)c1ccccc1 1 0 1D-0A 1 052.1631\,052.163 $15.885$
9CzFDPESPO O=P(c1ccccc1)(c1ccccc1)c1ccccc1Oc1ccc(-c2ccc(C3(c4ccc(-n5c6ccccc6c6ccccc65)cc4)c4ccccc4-c4ccccc43)cc2)cc1 1 0 1D-0A $851.986$ $14.245$
CTM O=C(c1ccc(-n2c3ccccc3c3ccccc32)cc1)c1ccc2sc3ccccc3c2c1 1 0 1D-0A $453.566$ $8.383$
tCTM CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(C(=O)c2ccc3sc4ccccc4c3c2)cc1 1 0 1D-0A $565.782$ $10.978$
PSBC CC(C)(C)c1ccc(-n2c(-c3ccc(-c4ccc(S(=O)(=O)c5ccc(-c6ccc(-n7c8ccccc8c8ccccc87)cc6)cc5)cc4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 1 0 1D-0A $884.117$ $15.560$
26DAcBPy CC1(C)c2ccccc2N(c2ccc(C(=O)c3cccc(C(=O)c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)n3)cc2)c2ccccc21 2 0 2D-0A $701.870$ $11.762$
25DAcBPy CC1(C)c2ccccc2N(c2ccc(C(=O)c3ccc(C(=O)c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)nc3)cc2)c2ccccc21 2 0 2D-0A $701.870$ $11.762$
26DPXZBPy O=C(c1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1cccc(C(=O)c2ccc(N3c4ccccc4Oc4ccccc43)cc2)n1 4 0 Multi-D/A $649.706$ $10.695$
BTITrz N#Cc1cc(-c2nc(-c3ccccc3)nc(-c3ccccc3)n2)ccc1-n1c2ccccc2c2c3ccccc3sc21 0 3 Multi-D/A $555.666$ $9.056$
mTPA–PPI c1ccc(N(c2ccccc2)c2ccc(-c3cccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccccc4)c3)cc2)cc1 1 0 1D-0A $613.764$ $12.136$
PPI c1ccc(-c2nc3c4ccccc4c4ccccc4c3n2-c2ccccc2)cc1 0 0 Unknown $370.455$ $6.999$
CzTPA-m-Trz c1ccc(-c2ccc(N(c3ccc(-c4ccc5c(c4)c4ccccc4n5-c4ccccc4)cc3)c3cccc(-c4nc(-c5ccccc5)nc(-c5ccccc5)n4)c3)cc2)cc1 2 1 D-A-D $793.974$ $14.773$
TPA–2PPITPA c1ccc(N(c2ccc(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cc3)cc2)c2ccc(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cc3)cc2)cc1 3 0 Multi-D/A 1 316.6251\,316.625 $26.055$
[…]
2Cz2tCzBn CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1c(-n2c3ccccc3c3ccccc32)cc(-n2c3ccccc3c3ccccc32)c(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c1C#N 4 2 Multi-D/A $988.336$ $19.137$
2tCz2CzBn CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c(-n2c3ccccc3c3ccccc32)c(C#N)c1-n1c2ccccc2c2ccccc21 4 2 Multi-D/A $988.336$ $19.137$
PmPmP c1cncc(-c2nc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)cc(-c3cncnc3)n2)c1 2 2 2D-2A $492.542$ $7.238$
PLQYs CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cc2ccc3cc(-n4c5ccc(C(C)(C)C)cc5c5cc(C(C)(C)C)ccc54)cc4ccc(c1)c2c34 2 0 2D-0A $757.078$ $15.968$
PXZ-ND c1ccc2c(c1)Oc1ccccc1N2c1ccc2ccc(N3c4ccccc4Oc4ccccc43)nc2n1 2 0 2D-0A $492.538$ $8.781$
CPPyC c1cc(-c2cncc(-n3c4ccccc4c4ccccc43)c2)cc(-n2c3ccccc3c3ccccc32)c1 2 0 2D-0A $485.590$ $8.943$
DMAC-ND CC1(C)c2ccccc2N(c2ccc3ccc(N4c5ccccc5C(C)(C)c5ccccc54)nc3n2)c2ccccc21 0 0 Unknown $544.702$ $9.848$
PrDPhAc c1ccc(C2(c3ccccc3)c3ccccc3N(c3cncc(N4c5ccccc5C(c5ccccc5)(c5ccccc5)c5ccccc54)n3)c3ccccc32)cc1 0 0 Unknown $742.926$ $12.808$
2Mi2SB Cc1nc2ccccc2n1-c1cc(S(=O)(=O)c2ccc(C(C)(C)C)cc2)c(-n2c(C)nc3ccccc32)cc1S(=O)(=O)c1ccc(C(C)(C)C)cc1 0 0 Unknown $730.956$ $9.242$
BFDMAc-NAI CC(C)(C)c1ccc(N2C(=O)c3cccc4c(N5c6ccccc6C(C)(C)c6ccc7c(oc8ccccc87)c65)ccc(c34)C2=O)cc1 1 0 1D-0A $626.756$ $10.946$
HPIPXZ c1ccc(N(c2ccc(-c3ccc(N(c4ccccc4)c4cccc5ccccc45)cc3)cc2)c2cccc3ccccc23)cc1 2 0 2D-0A $588.754$ $12.600$
3BPy-mDTC CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cc(C(=O)c2cccnc2)cc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c1 2 0 2D-0A $738.032$ $13.697$
2BPy-mDTC CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cc(C(=O)c2ccccn2)cc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c1 2 0 2D-0A $738.032$ $13.697$
BP-mDTC CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cc(C(=O)c2ccccc2)cc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c1 2 0 2D-0A $737.044$ $14.302$
Cz2ICz c1ccc2c(c1)c1ccccc1n2-c1ccc2c(c1)c1cccc3c4cc(-n5c6ccccc6c6ccccc65)ccc4n2c13 4 0 Multi-D/A $571.683$ $11.031$
Ac-BPCN CC1(C)c2ccccc2N(c2ccc3nc4c5cc(C#N)ccc5c5ccc(C#N)cc5c4nc3c2)c2ccccc21 1 6 Multi-D/A $537.626$ $8.942$
5CzICz c1ccc2c(c1)c1ccccc1n2-c1ccc2c(c1)c1cccc3c4ccccc4n2c31 3 0 Multi-D/A $406.488$ $7.934$
DCPDAPM CC1(C)c2ccccc2N(c2ccc(C(=O)c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)cc2)c2ccccc21 3 0 Multi-D/A $719.888$ $13.221$
SPFS-PXZ O=S1(=O)c2ccc(N3c4ccccc4Oc4ccccc43)cc2C2(c3ccccc3-c3ccccc32)c2cc(N3c4ccccc4Oc4ccccc43)ccc21 4 0 Multi-D/A $742.856$ $12.347$
DPS-PXZ O=S(=O)(c1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 0 Multi-D/A $580.665$ $9.670$
TX-CzPh O=c1c2ccccc2sc2ccc(-c3ccc4c(c3)c3ccccc3n4-c3ccccc3)cc12 1 0 1D-0A $453.566$ $8.179$
bpipH Brc1cccc(-c2nc3c4cccnc4c4ncccc4c3[nH]2)n1 0 0 Unknown $376.217$ $4.484$
Tri-PXZ-PCN N#Cc1c(-c2ccc(N3c4ccccc4Oc4ccccc43)cc2)nc(-c2ccc(N3c4ccccc4Oc4ccccc43)cc2)c(C#N)c1-c1ccc(N2c3ccccc3Oc3ccccc32)cc1 6 2 Multi-D/A $900.998$ $16.552$
ACID–BPPZ CC1(C)c2ccccc2N(c2ccc3c(c2)c2ccccc2c2nc4c5cccnc5c5ncccc5c4nc32)c2cc3c4ccccc4n(-c4ccccc4)c3cc21 2 2 2D-2A $754.897$ $13.392$
C-BPyIA c1ccc(-n2c(-c3ccc(-c4c5ccccc5c(-c5ccc(-c6nc7c8cccc9ccc%10cccc(c%10c98)c7n6-c6ccccc6)cc5)c5ccccc45)cc3)nc3c4cccc5ccc6cccc(c6c54)c32)cc1 0 0 Unknown $963.156$ $19.134$
N-BPyIA c1ccc(-c2nc3c4cccc5ccc6cccc(c6c54)c3n2-c2ccc(-c3c4ccccc4c(-c4ccc(-n5c(-c6ccccc6)nc6c7cccc8ccc9cccc(c9c87)c65)cc4)c4ccccc34)cc2)cc1 0 0 Unknown $963.156$ $19.134$
2-TA-AN c1ccc2c(c1)Sc1ccc(-c3c4ccccc4cc4ccccc34)cc1S2 0 0 Unknown $392.548$ $8.276$
ANOP-MeFAC Cc1ccc2c(c1)C1(c3ccccc3-c3ccccc31)c1cc(C)ccc1N2c1ccc2c3c(cccc13)-c1nc3nonc3nc1-2 1 0 1D-0A $603.685$ $9.577$
ANOP-DMAC CC1(C)c2ccccc2N(c2ccc3c4c(cccc24)-c2nc4nonc4nc2-3)c2ccccc21 1 0 1D-0A $453.505$ $6.923$
4PyCNICz N#Cc1c(-c2ccccc2)nc(-c2ccccc2)nc1-c1ccc(-n2c3ccccc3c3ccc4c(c5ccccc5n4-c4ccccc4)c32)cc1 2 2 2D-2A $663.784$ $11.543$
2PyCNBCz N#Cc1c(-c2ccccc2)nc(-c2ccc(-n3c4ccccc4c4cc(-c5ccc6c(c5)c5ccccc5n6-c5ccccc5)ccc43)cc2)nc1-c1ccccc1 2 2 2D-2A $739.882$ $13.210$
CzA-PPO Cc1ccc(N(c2ccc(C)cc2)c2c3ccccc3c(-c3ccc(P(=O)(c4ccccc4)c4ccc(-n5c6ccccc6c6ccccc65)cc4)cc3)c3ccccc23)cc1 2 0 2D-0A $814.969$ $14.483$
2PyCNICz N#Cc1c(-c2ccccc2)nc(-c2ccc(-n3c4ccccc4c4ccc5c(c6ccccc6n5-c5ccccc5)c43)cc2)nc1-c1ccccc1 2 2 2D-2A $663.784$ $11.543$
TAA-PPO Cc1ccc(N(c2ccc(C)cc2)c2c3ccccc3c(-c3ccc(P(=O)(c4ccccc4)c4ccccc4)cc3)c3ccccc23)cc1 1 0 1D-0A $649.774$ $11.386$
pSFIAc2 Cc1ccc2c(c1)c1c3c4c(c5c1n2-c1ccc(C(C)(C)C)cc1C51c2ccccc2-c2ccccc21)c1cc(C)ccc1n4-c1ccc(C(C)(C)C)cc1C31c2ccccc2-c2ccccc21 2 0 2D-0A $873.156$ $16.446$
TBN-TPA CC(C)(C)c1ccc(N2c3ccc(C(C)(C)C)cc3B3c4cc(C(C)(C)C)ccc4N(c4ccc(C(C)(C)C)cc4)c4cc(-n5c6ccc(C(C)(C)C)cc6c6cc(C(C)(C)C)ccc65)cc2c43)cc1 3 0 Multi-D/A $922.167$ $16.651$
CzBPDCb c1cc(-c2cccc(-n3c4ccccc4c4ncccc43)c2)cc(-n2c3ccccc3c3ccccc32)c1 1 0 1D-0A $485.590$ $8.943$
[2,1-b]IF C1=CC2(C=C3C1=CC1=C3C3(C4=c5ccccc5=CC4=C1)c1ccccc1Sc1ccccc13)c1ccccc1Sc1ccccc12 0 0 Unknown $618.826$ $9.206$
cyclohexanone BrCCCCCCC1(CCCCCCBr)c2cc(Br)ccc2-c2ccc(Br)cc21 0 0 Unknown $650.131$ $10.169$
thio-ether C1=CC2(C=C3C1=CC1=C3C3(C4=c5ccccc5=CC4=C1)c1ccccc1Sc1ccccc13)c1ccccc1Sc1ccccc12 0 0 Unknown $618.826$ $9.206$
ICz-2 c1ccc(-c2cccnc2-c2cc(-c3c[nH]nn3)cc(-c3ccccn3)n2)cc1 0 0 Unknown $376.423$ $4.658$
DDCzTrz c1ccc(-c2nc(-c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)nc(-c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)n2)cc1 4 1 Multi-D/A $970.152$ $17.261$
B-O-dpa CC(C)(C)c1cc2c3c(c1)N(c1ccccc1)c1ccccc1B3c1ccccc1O2 1 0 1D-0A $401.318$ $5.389$
B-dpa-SpiroAC CC(C)(C)c1cc2c3c(c1)N1c4ccccc4C4(c5ccccc5-c5ccccc54)c4cccc(c41)B3c1ccccc1N2c1ccccc1 2 0 2D-0A $638.623$ $9.743$
DCz-OTP c1cc(-c2ccccc2-c2cccc(-n3c4ccccc4c4ccccc43)c2)cc(-n2c3ccccc3c3ccccc32)c1 2 0 2D-0A $560.700$ $11.215$
B-dpa-Cz CC(C)(C)c1cc2c3c(c1)-n1c4ccccc4c4cccc(c41)B3c1ccccc1N2c1ccccc1 2 0 2D-0A $474.416$ $6.694$
c-D2T CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(N2c3ccccc3C3(c4ccccc4-c4cc5c(cc43)C3(c4ccccc4-5)c4ccccc4N(c4ccc(-c5nc(-c6ccccc6)nc(C6CCCCC6)n5)cc4)c4ccccc43)c3ccccc32)cc1 3 1 Multi-D/A 1 251.6351\,251.635 $23.238$
DPXZ-DPPM O=C(c1ccccc1)c1ccc2nc3c4ccc(N5c6ccccc6Oc6ccccc65)cc4c4cc(N5c6ccccc6Oc6ccccc65)ccc4c3nc2c1 4 2 Multi-D/A $746.826$ $13.471$
CBPMCN N#Cc1ccc(-n2c(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)nc3c4ccccc4c4ccccc4c32)cc1 1 2 A-D-A $560.660$ $9.968$
CPBPMCN N#Cc1ccc(-n2c(-c3ccc(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 1 2 A-D-A $636.758$ $11.635$
TPB-AC N#Cc1ccc(-c2cc(-c3ccccc3)c(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)cc2-c2ccccc2)cc1 1 2 A-D-A $574.727$ $11.696$
Me-DMAC Cc1nc(C)c(C#N)c(-c2ccc(N3c4ccccc4C(C)(C)c4ccccc43)cc2)c1C#N 1 2 A-D-A $440.550$ $7.218$
iTPBI-CN N#Cc1c(-n2c(-c3ccccc3)nc3ccccc32)cc(-n2c(-c3ccccc3)nc3ccccc32)cc1-n1c(-c2ccccc2)nc2ccccc21 0 2 0D-2A $679.787$ $10.576$
TCP c1ccc2c(c1)c1ccccc1n2-c1cc(-n2c3ccccc3c3ccccc32)cc(-n2c3ccccc3c3ccccc32)c1 3 0 Multi-D/A $573.699$ $10.978$
CDBP-BP-PXZ Cc1cc(-n2c3ccccc3c3ccccc32)ccc1-c1ccc(-n2c3ccccc3c3cc(C(=O)c4ccc(N5c6ccccc6Oc6ccccc65)cc4)ccc32)cc1C 4 0 Multi-D/A $797.958$ $14.971$
CDBP-BP-DMAC Cc1cc(-n2c3ccccc3c3ccccc32)ccc1-c1ccc(-n2c3ccccc3c3cc(C(=O)c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)ccc32)cc1C 3 0 Multi-D/A $824.040$ $15.505$
3CzPhpPM c1ccc(-c2ccc3c(c2)c2ccccc2n3-c2ccc(-c3cc(-c4ccccc4)nc(-c4ccccc4)n3)cc2)cc1 1 1 D-A $549.677$ $10.242$
SFI23pPM c1ccc(-c2cc(-c3ccc(-n4c5ccccc5c5cc6c(cc54)C4(c5ccccc5-c5ccccc54)c4ccccc4-6)cc3)nc(-c3ccccc3)n2)cc1 1 1 D-A $711.868$ $12.918$
3PCzPFP c1ccc(-n2c3ccccc3c3cc(-c4cnc5oc6ncccc6c5c4)ccc32)cc1 1 0 1D-0A $411.464$ $7.140$
SFI34pPM c1ccc(-c2cc(-c3ccc(-n4c5ccccc5c5c6c(ccc54)-c4ccccc4C64c5ccccc5-c5ccccc54)cc3)nc(-c3ccccc3)n2)cc1 1 1 D-A $711.868$ $12.918$
SBA-2DPS O=S(=O)(c1ccccc1)c1ccc(N2c3ccccc3C3(c4ccccc42)c2ccccc2N(c2ccc(S(=O)(=O)c4ccccc4)cc2)c2ccccc23)cc1 2 0 2D-0A $778.955$ $11.301$
2DMAC-BP-F CC1(C)c2ccccc2N(c2ccc3c(c2)c2cc(N4c5ccccc5C(C)(C)c5ccccc54)ccc2c2nc4c(F)cccc4nc32)c2ccccc21 2 2 2D-2A $712.872$ $13.447$
DCA N#Cc1cc(C#N)cc(-c2c3ccccc3c(-c3cc(C#N)cc(C#N)c3)c3ccccc23)c1 0 8 Multi-D/A $430.470$ $6.814$
DPFPhe CCCCCCN1C2=CCC(=CC3c4cc(-c5ccccc5)ccc4-c4ccc(-c5ccccc5)cc43)C=C2Oc2ccccc21 0 0 Unknown $597.802$ $11.710$
DFPFPhe CCCCCCN1C2=CCC(=CC3c4cc(-c5ccc(F)cc5)ccc4-c4ccc(-c5ccc(F)cc5)cc43)C=C2Oc2ccccc21 0 0 Unknown $633.782$ $11.988$
3ai CCCCc1cc2c(s1)-c1sc(CCCC)cc1P2(=O)c1ccc(-n2c3ccccc3c3ccccc32)cc1 1 0 1D-0A $565.744$ $9.212$
DNFPh CCCCCCN1C2=CCC(=CC3c4cc(-c5cccc6ccccc56)ccc4-c4ccc(-c5cccc6ccccc56)cc43)C=C2Sc2ccccc21 0 0 Unknown $713.990$ $14.730$
pCzAnBzt N#Cc1ccc(-c2c3ccccc3c(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)c3ccccc23)cc1 1 2 A-D-A $520.635$ $10.296$
DPE-DPXZ C(#Cc1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 0 Multi-D/A $540.622$ $10.237$
3CzAnBzt N#Cc1ccc(-c2c3ccccc3c(-c3ccc4c(c3)c3ccccc3n4-c3ccccc3)c3ccccc23)cc1 1 2 A-D-A $520.635$ $10.296$
[…]
ADN c1ccc2cc(-c3c4ccccc4c(-c4ccc5ccccc5c4)c4ccccc34)ccc2c1 0 0 Unknown $430.550$ $9.633$
Et3SiH FC1=C(F)C(c2c(F)c(F)c(F)c(F)c2F)(c2c(F)c(F)c(F)c(F)c2F)C(c2c(F)c(F)c(F)c(F)c2F)(c2c(F)c(F)c(F)c(F)c2F)B1c1c(F)c(F)c(F)c(F)c1F 0 0 Unknown $932.132$ $11.081$
DPE-DDMAc CC1(C)c2ccccc2N(c2ccc(C#Cc3ccc(N4c5ccccc5C(C)(C)c5ccccc54)cc3)cc2)c2ccccc21 2 0 2D-0A $592.786$ $11.305$
mCz-TAn-CN CC(C)(C)c1ccc2c(-c3cccc(-n4c5ccccc5c5ccccc54)c3)c3cc(C(C)(C)C)ccc3c(-c3ccc(C#N)cc3)c2c1 1 2 A-D-A $632.851$ $12.891$
indenoacridine c1ccc2[nH]c3ccc4c5ccccc5cc4c3cc2c1 1 0 1D-0A $267.331$ $5.628$
MPAc Cc1ccc2c(c1)C(c1ccccc1)(c1ccccc1)c1cc(C)ccc1N2 0 0 Unknown $361.488$ $6.743$
TBAc CC(C)(C)C1=CC=CC(C(C)(C)C)(C2(C3(C(C)(C)C)C=CC=C(C(C)(C)C)C3)c3ccccc3N(c3ccccc3)c3ccccc32)C1 1 0 1D-0A $637.996$ $13.686$
poCz-SO O=P(c1ccccc1)(c1ccccc1)c1ccc2[nH]c3ccc(P(=O)(c4ccccc4)c4ccccc4)cc3c2c1 1 0 1D-0A $567.565$ $6.600$
H4TCPE O=C(O)c1ccc(CC(c2ccc(C(=O)O)cc2)(c2ccc(C(=O)O)cc2)c2ccc(C(=O)O)cc2)cc1 0 0 Unknown $510.498$ $5.057$
CbBPCb c1cc(-c2cccc(-n3c4ccccc4c4cccnc43)c2)cc(-n2c3ccccc3c3cccnc32)c1 0 0 Unknown $486.578$ $8.338$
TMOE COc1ccc(C(=C(c2ccc(OC)cc2)c2ccc(OC)cc2)c2ccc(OC)cc2)cc1 0 0 Unknown $452.550$ $6.728$
FAPbCl3 c1ccc(-c2ccc(-c3ccccc3)c(-c3ccccc3)c2-c2ccccc2)cc1 0 0 Unknown $382.506$ $8.355$
DBT-SADF O=S1(=O)c2ccccc2-c2cc(N3c4ccccc4C4(c5ccccc5-c5ccccc54)c4ccccc43)ccc21 1 0 1D-0A $545.663$ $8.646$
mPXTMP Cc1c(-c2ccc3c(c2)Oc2ccccc2S3)cccc1-c1ccc2c(c1)Oc1ccccc1S2 0 0 Unknown $488.633$ $9.843$
APDC-tPh N#Cc1nc2c(nc1C#N)-c1ccc(-c3cc(-c4ccccc4)cc(-c4ccccc4)c3)c3cccc-2c13 0 2 0D-2A $482.546$ $8.022$
POZDBPHZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2c(ccc3nc4ccc5cc(N6c7ccccc7Oc7ccccc76)ccc5c4nc32)c1 4 2 Multi-D/A $642.718$ $12.240$
CPTBF CC(C)(C)c1ccc2c(c1)C(c1ccc(-n3c4ccccc4c4ccccc43)cc1)(c1ccc(-n3c4ccccc4c4ccccc43)cc1)c1cc(C(C)(C)C)ccc1-2 2 0 2D-0A $761.025$ $14.839$
trimesitylboron Cc1cc(C)c(B(c2c(C)cc(C)cc2C)c2c(C)cc(C)cc2C)c(C)c1 0 0 Unknown $368.373$ $4.979$
CMP c1ccc(C(c2ccccc2)(c2ccccc2)c2ccccc2)cc1 0 0 Unknown $320.435$ $6.069$
CH3OH–CHCl3 O=P(c1ccccc1)(c1ccccc1)c1cccc(-c2ccc3ccc4cccc5ccc2c3c45)c1 0 0 Unknown $478.531$ $7.890$
s-triazine CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc([Si](c2ccccc2)(c2ccccc2)c2ccc(B(O)O)cc2)cc1 1 0 1D-0A $657.739$ $6.436$
[BMIM][PF6] CCCCCCC1(CCCCCC)c2ccccc2-c2ccc(Br)cc21 0 0 Unknown $413.443$ $8.656$
1SBFN c1ccc2c(c1)-c1ccccc1C21c2ccccc2-c2ccc(-c3cccc4ccccc34)cc21 0 0 Unknown $442.561$ $8.850$
2SBFN c1ccc2c(c1)-c1ccccc1C21c2ccccc2-c2ccc(-c3ccc4ccccc4c3)cc21 0 0 Unknown $442.561$ $8.850$
F2PA CC1(C)c2ccccc2-c2ccc(N(c3ccccc3)c3ccc4c(c3)C(C)(C)c3ccccc3-4)cc21 1 0 1D-0A $477.651$ $9.769$
TSBFB c1ccc2c(c1)-c1ccccc1C21c2ccccc2-c2ccc(-c3cc(-c4ccc5c(c4)C4(c6ccccc6-c6ccccc64)c4ccccc4-5)cc(-c4ccc5c(c4)C4(c6ccccc6-c6ccccc64)c4ccccc4-5)c3)cc21 0 0 Unknown 1 021.2751\,021.275 $19.718$
DPPP c1ccc(-c2cc(-c3ccccc3)cc(-c3ccccc3-c3ccc4ccc5c(-c6ccccc6-c6cc(-c7ccccc7)cc(-c7ccccc7)c6)ccc6ccc3c4c65)c2)cc1 0 0 Unknown $811.040$ $17.920$
TPA-NZP c1ccc(-c2c3ccccc3c(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)c3nsnc23)cc1 1 0 1D-0A $505.646$ $9.648$
mTPA-PPI c1ccc(N(c2ccccc2)c2ccc(-c3cccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccccc4)c3)cc2)cc1 1 0 1D-0A $613.764$ $12.136$
BSBFB c1cc(-c2ccc3c(c2)C2(c4ccccc4-c4ccccc42)c2ccccc2-3)cc(-c2ccc3c(c2)C2(c4ccccc4-c4ccccc42)c2ccccc2-3)c1 0 0 Unknown $706.888$ $13.708$
BN-DBC-Ph2 C1=CB2C(=C3C(=C4C=C(c5ccccc5)C=CN4c4ccc(-c5ccccc5)cc43)c3ccccc32)C=C1 0 0 Unknown $481.407$ $7.869$
TCbTrz c1ccc(-c2nc(-c3ccccc3)nc(-c3cc(-n4c5ccccc5c5ncccc54)c(-n4c5ccccc5c5ncccc54)c(-n4c5ccccc5c5ncccc54)c3)n2)cc1 0 1 0D-1A $807.921$ $12.349$
Xm-mCP O=c1c2ccccc2oc2ccc(-c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)cc12 2 0 2D-0A $602.693$ $10.807$
Xp-mCP CC1(C)OB(c2cc(-n3c4ccccc4c4ccccc43)cc(-n3c4ccccc4c4ccccc43)c2)OC1(C)C 2 0 2D-0A $534.468$ $8.180$
ethylacetate O=S(=O)(c1ccc(N2c3ccccc3C=Cc3ccccc32)cc1)c1ccc(N2c3ccccc3C=Cc3ccccc32)cc1 2 0 2D-0A $600.743$ $10.427$
3DMACIPN CC1(C)c2ccccc2N(c2c(F)c(N3c4ccccc4C(C)(C)c4ccccc43)c(C#N)c(N3c4ccccc4C(C)(C)c4ccccc43)c2C#N)c2ccccc21 3 4 Multi-D/A $767.952$ $13.896$
ICZ CCCCCc1c2[nH]c3ccccc3c2cc2[nH]c3ccccc3c12 2 0 2D-0A $326.443$ $6.688$
BDAVBi C(=Cc1ccc(N(c2ccccc2)c2ccccc2)cc1)c1ccc(-c2ccc(C=Cc3ccc(N(c4ccccc4)c4ccccc4)cc3)cc2)cc1 2 0 2D-0A $692.906$ $14.634$
DPVBi C(=C(c1ccccc1)c1ccccc1)c1ccc(-c2ccc(C=C(c3ccccc3)c3ccccc3)cc2)cc1 0 0 Unknown $510.680$ $10.531$
CBTD2 CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-c2ccc(-c3ccc(-n4c5ccc(C(C)(C)C)cc5c5cc(C(C)(C)C)ccc54)cc3)c3nsnc23)cc1 2 0 2D-0A $843.197$ $16.409$
MeOCzPCN COc1ccc2c(c1)c1cc(OC)ccc1n2-c1ncnc(-n2c3ccc(OC)cc3c3cc(OC)ccc32)c1C#N 2 2 2D-2A $555.594$ $6.577$
Na-DMAC CC1(C)c2ccccc2N(c2ccc(-c3c(C#N)c(-c4cccc5ccccc45)nc(-c4cccc5ccccc45)c3C#N)cc2)c2ccccc21 1 2 A-D-A $664.812$ $12.241$
3Py-DMAC CC1(C)c2ccccc2N(c2ccc(-c3c(C#N)c(-c4cccnc4)nc(-c4cccnc4)c3C#N)cc2)c2ccccc21 1 2 A-D-A $566.668$ $8.725$
6tBPA CC(C)(C)c1ccc(N(c2ccc(C(C)(C)C)cc2)c2c3ccc(C(C)(C)C)cc3c(N(c3ccc(C(C)(C)C)cc3)c3ccc(C(C)(C)C)cc3)c3ccc(C(C)(C)C)cc23)cc1 2 0 2D-0A $849.304$ $18.718$
C-O COc1ccc2c(c1)c1ccccc1n2-c1cc2c(cc1-n1c3ccccc3c3ccccc31)Cc1ccccc1O2 2 0 2D-0A $542.638$ $9.586$
IDFL-2DPA CC1(C)c2ccccc2-c2c1cc1c3cc(N(c4ccccc4)c4ccccc4)ccc3n3c4ccc(N(c5ccccc5)c5ccccc5)cc4c2c13 4 0 Multi-D/A $691.878$ $14.083$
ACRF N#Cc1ccc2c(c1)C1(c3cc(C#N)ccc3-2)c2ccccc2N(c2ccccc2)c2ccccc21 1 4 Multi-D/A $457.536$ $7.576$
TPSi-F c1ccc(C2(c3ccc([Si](c4ccccc4)(c4ccccc4)c4ccccc4)cc3)c3ccccc3-c3ccccc32)cc1 0 0 Unknown $576.815$ $7.427$
PAB c1ccc(N(c2ccccc2)c2cc3c4c(c2)N(c2ccccc2)c2ccccc2B4c2ccccc2N3c2ccccc2)cc1 3 0 Multi-D/A $587.535$ $9.239$
2tPAB CC(C)(C)c1ccc(N2c3ccc(C(C)(C)C)cc3B3c4cc(C(C)(C)C)ccc4N(c4ccc(C(C)(C)C)cc4)c4cc(N(c5ccccc5)c5ccccc5)cc2c43)cc1 3 0 Multi-D/A $811.967$ $14.429$
pDTPA-DTPZ c1ccc(N(c2ccccc2)c2ccc(-c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)c4nc5c(nc34)c3sccc3c3ccsc35)cc2)cc1 2 2 2D-2A $779.006$ $15.486$
oDTPA-DTPZ c1ccc(N(c2ccccc2)c2ccc(-c3cc4nc5c(nc4cc3-c3ccc(N(c4ccccc4)c4ccccc4)cc3)c3sccc3c3ccsc35)cc2)cc1 2 2 2D-2A $779.006$ $15.486$
oTPAPO-DTPZ O=P(c1ccccc1)(c1ccccc1)c1cc2nc3c(nc2cc1-c1ccc(N(c2ccccc2)c2ccccc2)cc1)c1sccc1c1ccsc13 1 2 A-D-A $735.874$ $11.989$
BPZPP c1ccc(-c2cc3nc4c5ccc(N6c7ccccc7Oc7ccccc76)nc5c5nc(N6c7ccccc7Oc7ccccc76)ccc5c4nc3cc2-c2ccccc2)cc1 2 2 2D-2A $796.890$ $14.364$
BIPOCz c1ccc(-c2nc(-c3ccc(OCCCCc4cccc5c4[nH]c4ccccc45)cc3)n(-c3ccccc3)c2-c2ccccc2)cc1 1 0 1D-0A $609.773$ $10.909$
tDPAC-BN CC(C)(C)c1ccc(N2c3ccc(C(C)(C)C)cc3B3c4cc(C(C)(C)C)ccc4N(c4ccc(C(C)(C)C)cc4)c4cc(N5c6ccccc6C(c6ccccc6)(c6ccccc6)c6ccccc65)cc2c43)cc1 3 0 Multi-D/A $976.174$ $17.125$
tDMAC-BN CC(C)(C)c1ccc(N(c2ccc(C(C)(C)C)cc2)c2cc3c4c(c2)N2c5ccccc5C(C)(C)c5cccc(c52)B4c2cc(C(C)(C)C)ccc2N3c2ccc(C(C)(C)C)cc2)cc1 3 0 Multi-D/A $852.032$ $15.068$
2ICz-OXD c1ccc(-n2c3ccccc3c3c2ccc2c4ccccc4n(-c4ccc(-c5nnc(-c6ccc(-n7c8ccccc8c8ccc9c(c%10ccccc%10n9-c9ccccc9)c87)cc6)o5)cc4)c23)cc1 4 0 Multi-D/A $883.027$ $15.792$
SIR-B Cc1ccc2c(c1)C1(c3ccc(-c4ccccc4)cc3C(=O)c3ccc(-c4ccccc4)cc31)c1cc(C)ccc1N2c1ccc(C(C)(C)C)cc1 1 0 1D-0A $671.884$ $12.645$
SIR-C Cc1ccc2c(c1)C1(c3ccc(-c4ccccc4-c4ccccc4)cc3C(=O)c3ccc(-c4ccccc4-c4ccccc4)cc31)c1cc(C)ccc1N2c1ccc(C(C)(C)C)cc1 1 0 1D-0A $824.080$ $15.979$
SIR-A Cc1ccc2c(c1)C1(c3ccccc3C(=O)c3ccccc31)c1cc(C)ccc1N2c1ccc(C(C)(C)C)cc1 1 0 1D-0A $519.688$ $9.311$
TRZ3DPA c1ccc(N(c2ccccc2)c2ccc(-c3nc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)nc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)n3)cc2)cc1 3 1 Multi-D/A $811.005$ $15.282$
BO3N Nc1ccc2c(c1)Oc1cc(N)cc3c1B2c1ccc(Nc2cccc4cc5ccc6cc7cc8ccccc8cc7cc6c5cc24)cc1O3 0 0 Unknown $641.539$ $9.088$
BPPyA Cc1ccc(C)c(N(c2ccc3ccc4c(N(c5cc(C)ccc5C)c5cccc6c5oc5ccccc56)ccc5ccc2c3c54)c2cccc3c2oc2ccccc23)c1 2 0 2D-0A $772.948$ $16.556$
SiCN2Cz N#Cc1cc(-c2nc(-n3c4ccccc4c4ccccc43)nc(-n3c4ccccc4c4ccccc43)n2)cc([Si](c2ccccc2)(c2ccccc2)c2ccccc2)c1 2 3 Multi-D/A $770.972$ $8.982$
PhPC c1ccc(-c2c3ccccc3c(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)c3ccccc23)cc1 1 0 1D-0A $495.625$ $10.424$
1MOBI COc1cc(/C=C/c2ccc(-n3cnc4ccccc43)cc2)c(OC)cc1/C=C/c1ccc(-n2cnc3ccccc32)cc1 0 0 Unknown $574.684$ $8.722$
3BrBI Brc1cc(/C=C/c2ccc(-n3cnc4ccccc43)cc2)c(Br)cc1/C=C/c1ccc(-n2cnc3ccccc32)cc1 0 0 Unknown $672.424$ $10.230$
DMBFTX CC1(C)c2ccccc2-c2cccc(-c3ccc4sc5ccc(-c6cccc7c6C(C)(C)c6ccccc6-7)cc5c(=O)c4c3)c21 0 0 Unknown $596.795$ $11.361$
SBF-AnCN N#Cc1ccc(-c2c3ccccc3c(-c3ccc4c(c3)C3(c5ccccc5-c5ccccc53)c3ccccc3-4)c3ccccc23)cc1 0 2 0D-2A $593.729$ $11.542$
DMF-AnCN CC1(C)c2ccccc2-c2ccc(-c3c4ccccc4c(-c4ccc(C#N)cc4)c4ccccc34)cc21 0 2 0D-2A $471.603$ $9.505$
DPF-AnCN N#Cc1ccc(-c2c3ccccc3c(-c3ccc4c(c3)C(c3ccccc3)(c3ccccc3)c3ccccc3-4)c3ccccc23)cc1 0 2 0D-2A $595.745$ $11.562$
DPA c1ccc(-c2c3ccccc3c(-c3ccccc3)c3ccccc23)cc1 0 0 Unknown $330.430$ $7.327$
TPA–DAP–TPA c1ccc(N(c2ccccc2)c2ccc(-c3c4ccccc4c(-c4ccc5ccc6c(-c7c8ccccc8c(-c8ccc(N(c9ccccc9)c9ccccc9)cc8)c8ccccc78)ccc7ccc4c5c76)c4ccccc34)cc2)cc1 2 0 2D-0A 1 041.3101\,041.310 $22.804$
AcDbp CC1(C)c2ccccc2N(c2ccc3c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4c4nc5ccccc5nc4c3c2)c2ccccc21 2 2 2D-2A $694.882$ $13.308$
4-Ac CC(C)(C)c1ccc2c(c1)C(C)(C)c1cc(C(C)(C)C)ccc1N2c1ccc(C(F)(F)F)cc1 1 0 1D-0A $465.603$ $9.409$
czpzpy c1cc(-c2cccc3c2[nH]c2ccccc23)nc(-n2cccn2)c1 1 0 1D-0A $310.360$ $4.569$
NyDPO c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2ccc3nc(-c4ccc(N5c6ccccc6Oc6ccccc65)cc4)ccc3n2)cc1 4 0 Multi-D/A $644.734$ $12.115$
1MPA Cc1nc(N2c3ccccc3C(c3ccccc3)(c3ccccc3)c3ccccc32)cc(N2c3ccccc3C(c3ccccc3)(c3ccccc3)c3ccccc32)n1 0 1 0D-1A $756.953$ $13.117$
LEt CCc1ccccc1P(c1ccccc1CC)c1ccccc1P(c1ccccc1CC)c1ccccc1CC 0 0 Unknown $558.686$ $7.453$
DC-TC CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(C(=O)C(=O)c2ccc(-n3c4ccc(C(C)(C)C)cc4c4cc(C(C)(C)C)ccc43)cc2)cc1 2 0 2D-0A $765.054$ $14.136$
DC-ACR CC1(C)c2ccccc2N(c2ccc(C(=O)C(=O)c3ccc(N4c5ccccc5C(C)(C)c5ccccc54)cc3)cc2)c2ccccc21 2 0 2D-0A $624.784$ $10.970$
1G-F1 c1ccc(P(c2ccccc2)c2cscc2P(c2ccccc2)c2ccccc2)cc1 0 0 Unknown $452.499$ $5.265$
P–Cu–C c1ccc(P(c2ccccc2)c2ccccc2-c2ccccc2P(c2ccccc2)c2ccccc2)cc1 0 0 Unknown $522.568$ $6.870$
BPCP-2CPC N#Cc1cc(-n2c3ccccc3c3cc(-c4ccc(C5(c6ccc(-n7c8ccccc8c8ccccc87)cc6)CCCCC5)cc4)ccc32)c(-n2c3ccccc3c3cc(-c4ccc(C5(c6ccc(-n7c8ccccc8c8ccccc87)cc6)CCCCC5)cc4)ccc32)cc1C#N 4 4 Multi-D/A 1 257.5981\,257.598 $23.653$
2H-Qz c1ccc(-c2cc(-c3ccccn3)cc(-c3nc(-c4ccc(N5c6ccccc6Oc6ccccc65)cc4)c4ccccc4n3)c2)nc1 2 1 D-A-D $617.712$ $10.663$
PPINCN-Cz N#Cc1ccc(-n2c(-c3ccc4ccc5cccc6ccc3c4c56)nc3c4ccc(-c5ccc(-n6c7ccccc7c7ccccc76)cc5)cc4c4cc(-c5ccc(-n6c7ccccc7c7ccccc76)cc5)ccc4c32)c2ccccc12 2 2 2D-2A 1 052.2531\,052.253 $20.449$
PPINCN N#Cc1ccc(-n2c(-c3ccc4ccc5cccc6ccc3c4c56)nc3c4ccccc4c4ccccc4c32)c2ccccc12 0 2 0D-2A $569.667$ $10.921$
BDAPM O=C(c1ccc(N2c3ccccc3[Si](c3ccccc3)(c3ccccc3)c3ccccc32)cc1)c1ccc(N2c3ccccc3[Si](c3ccccc3)(c3ccccc3)c3ccccc32)cc1 2 0 2D-0A $877.208$ $9.239$
12AcCz-PM CC1(C)c2ccccc2N(c2ccc(-c3cc(-c4ccccc4)nc(-c4ccccc4)n3)cc2)c2c1ccc1c2c2ccccc2n1-c1ccccc1 2 1 D-A-D $680.855$ $12.684$
23AcCz-PM CC1(C)c2ccccc2N(c2ccc(-c3cc(-c4ccccc4)nc(-c4ccccc4)n3)cc2)c2cc3c(cc21)c1ccccc1n3-c1ccccc1 2 1 D-A-D $680.855$ $12.684$
CDDPI c1ccc2c(-n3c4ccccc4c4ccccc43)ccc(-c3nc4c5ccccc5c5ccccc5c4n3-c3ccc4c(c3)OCCO4)c2c1 1 0 1D-0A $643.746$ $11.020$
DDBP c1ccc(-c2nc(-c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)c4ccccc34)n(-c3ccc4c(c3)OCCO4)c2-c2ccccc2)cc1 1 0 1D-0A $723.876$ $12.934$
CCDD O=C(c1ccccc1)c1ccc(-n2c3ccc(N(c4ccccc4)c4ccccc4)cc3c3cc(N(c4ccccc4)c4ccccc4)ccc32)cc1 3 0 Multi-D/A $681.839$ $12.954$
DPNC COc1ccc(N(c2ccc(OC)cc2)c2ccc3c(c2)c2cc(N(c4ccc(OC)cc4)c4ccc(OC)cc4)ccc2n3-c2cccc3ccccc23)cc1 3 0 Multi-D/A $747.895$ $12.911$
LiPr CC(C)c1ccccc1P(c1ccccc1C(C)C)c1ccccc1P(c1ccccc1C(C)C)c1ccccc1C(C)C 0 0 Unknown $614.794$ $9.697$
Spiro-CN Cc1ccc(N(c2ccc(C)cc2)c2ccc3c(c2)C2(c4cc(C#N)ccc4-c4ccc(C#N)cc42)c2cc(N(c4ccc(C)cc4)c4ccc(C)cc4)ccc2-3)cc1 2 4 Multi-D/A $756.953$ $13.947$
TMDBP CC1(C)c2ccccc2Nc2c1ccc1c2C(C)(C)c2ccccc2N1 0 0 Unknown $340.470$ $6.452$
ACRSA O=C1c2ccccc2C2(c3ccccc31)c1ccccc1N(c1ccccc1)c1ccccc12 1 0 1D-0A $435.526$ $7.397$
CAPI c1ccc2c(-n3c(-c4ccc(-c5c6ccccc6c(-n6c7ccccc7c7ccccc76)c6ccccc56)cc4)nc4c5ccccc5c5ccccc5c43)cccc2c1 1 0 1D-0A $761.928$ $15.223$
PhCzSPOTz O=P(c1ccccc1)(c1ccccc1)c1nc(-c2ccccc2)nc(-n2c3ccccc3c3ccccc32)n1 1 1 D-A $522.548$ $6.275$
CCAPI N#Cc1ccc(-n2c(-c3ccc(-c4c5ccccc5c(-n5c6ccccc6c6ccccc65)c5ccccc45)cc3)nc3c4ccccc4c4ccccc4c32)c2ccccc12 1 2 A-D-A $786.938$ $15.094$
2CzCbPy c1cc(-n2c3ccccc3c3ccccc32)nc(-n2c3ccc(-n4c5ccccc5c5ccccc54)cc3c3ncccc32)c1 2 0 2D-0A $575.675$ $9.768$
CzCbPy c1ccc(-n2c3ccc(-n4c5ccccc5c5ccccc54)cc3c3ncccc32)nc1 1 0 1D-0A $410.480$ $6.671$
DCE CC(C)CCC[C@@H](C)[C@H]1CC[C@H]2[C@@H]3CC=C4C[C@@H](O)CC[C@]4(C)[C@H]3CC[C@]12C 0 0 Unknown $386.664$ $7.389$
PICFOCz FC(F)(F)c1cccc(-n2c(-c3ccc(OCCCCn4c5ccccc5c5ccccc54)cc3)nc3c4ccccc4c4ccccc4c32)c1 1 0 1D-0A $675.754$ $11.985$
dpmt Cc1sc(C)c(P(c2ccccc2)c2ccccc2)c1P(c1ccccc1)c1ccccc1 0 0 Unknown $480.553$ $5.881$
BICFOCz FC(F)(F)c1cccc(-n2c(-c3ccc(OCCCCn4c5ccccc5c5ccccc54)cc3)nc(-c3ccccc3)c2-c2ccccc2)c1 1 0 1D-0A $677.770$ $11.859$
19a c1ccc2cc(-c3ccc4cc5cc(-c6ccc7ccccc7c6)ccc5cc4c3)ccc2c1 0 0 Unknown $430.550$ $9.633$
BTPBFCz c1ccc(-c2nc(-c3ccccc3-n3c4ccccc4c4c5oc6ccccc6c5ccc43)c3sc4ccccc4c3n2)cc1 1 1 D-A $593.711$ $11.175$
Ph-OBNA c1ccc(-n2c3ccccc3c3c4c5c(cc32)Oc2ccccc2B5c2ccccc2O4)cc1 1 0 1D-0A $435.291$ $5.511$
HPVAnt CCCCCCc1ccc(C=Cc2ccc3cc4ccccc4cc3c2)cc1 0 0 Unknown $364.532$ $8.286$
BTPBFCz-D3 c1ccc(-c2nc(-c3ccc(-n4c5ccccc5c5ccc6c(c7ccccc7n6-c6ccccc6)c54)cc3-n3c4ccccc4c4c5oc6ccccc6c5ccc43)c3sc4ccccc4c3n2)cc1 3 1 Multi-D/A $924.101$ $17.369$
ICBNTrz4 CC1(C)c2ccccc2-c2cc3c4ccccc4n(-c4ccc(-c5nc(-c6ccccc6)nc(-c6ccccc6)n5)cc4C#N)c3cc21 1 3 Multi-D/A $615.740$ $10.148$
1PCTRz c1ccc(-c2nc(-c3ccccc3)nc(-c3ccc(-n4c5ccccc5c5cc(-n6c7ccccc7c7ccccc76)ccc54)c(-c4ccccc4)c3)n2)cc1 2 1 D-A-D $715.860$ $12.734$
DMACN-B Cc1cc(C)c(N(c2ccc3c(c2)C(c2ccccc2)(c2ccccc2)c2cccc4c2N3c2ccccc2C4(C)C)c2c(C)cc(C)cc2C)c(C)c1 2 0 2D-0A $700.970$ $13.812$
CzSO-1CO CC1(OCOC(=O)c2ccccc2)CC=Cc2c1c1ccccc1n2-c1ccc(S(=O)(=O)c2ccccc2)cc1 0 0 Unknown $549.648$ $6.927$
PPO21 O=P(c1ccccc1)(c1ccccc1)c1ccc(-n2c3ccccc3c3cc(P(=O)(c4ccccc4)c4ccccc4)ccc32)cc1 1 0 1D-0A $643.663$ $8.063$
OXD-7 CC(C)(C)c1ccc(-c2nnc(-c3cccc(-c4nnc(-c5ccc(C(C)(C)C)cc5)o4)c3)o2)cc1 0 0 Unknown $478.596$ $7.716$
TCZ1 CCCCC($C$)Cn1c2ccc(-n3c4ccccc4c4ccccc43)cc2c2cc(-n3c4ccccc4c4ccccc43)ccc21 3 0 Multi-D/A $609.817$ $12.205$
APPT c1cc2c3c(cccc3c1)-c1nc3c4cccnc4c4ncccc4c3nc1-2 0 1 0D-1A $356.388$ $5.527$
ANQDC N#Cc1cc2nc3c(nc2cc1C#N)-c1cccc2cccc-3c12 0 5 Multi-D/A $304.312$ $4.174$
ECDPTT CCn1c2ccccc2c2cc(C(=C(c3ccccc3)c3ccccc3)c3ccc(-c4ccc5c(c4)S(=O)(=O)c4ccccc4S5(=O)=O)cc3)ccc21 1 0 1D-0A $727.907$ $10.468$
1CM CCCCCCCCn1c2ccccc2c2cc(Nc3nc(OC)nc(OC)n3)ccc21 1 1 D-A $433.556$ $6.101$
NPE-AcDPS CC(c1ccccc1)N1c2cc(N3c4ccccc4C(C)(C)c4ccccc43)ccc2S(=O)(=O)c2ccc(N3c4ccccc4C(C)(C)c4ccccc43)cc21 3 0 Multi-D/A $749.980$ $12.950$
DMAC-Cz CC1(C)c2ccccc2N(c2ccc3[nH]c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4c3c2)c2ccccc21 3 0 Multi-D/A $581.763$ $11.539$
VBC-PTZ CCN($C$)c1ccc2c(c1)oc(=O)c1c(=O)oc3cc(N4c5ccccc5Sc5ccccc54)ccc3c12 2 0 2D-0A $532.621$ $7.833$
t-BuPCz CC(C)(C)c1ccc(-c2ccc3[nH]c4ccc(-c5ccc(C(C)(C)C)cc5)cc4c3c2)cc1 1 0 1D-0A $431.623$ $9.250$
4BPy-mDTC CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cc(C(=O)c2ccncc2)cc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c1 2 0 2D-0A $738.032$ $13.697$
DC–TC Cc1ccc(N(c2ccc(C)cc2)c2ccc(C3(c4ccc(N(c5ccc(C)cc5)c5ccc(C)cc5)cc4)CCCCC3)cc2)cc1 2 0 2D-0A $626.888$ $13.110$
DFDBQPXZ Fc1cc2nc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)nc2cc1F 4 1 Multi-D/A $680.714$ $12.393$
DPE C(#Cc1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 0 Multi-D/A $540.622$ $10.237$
CBZ CCCCCCOc1ccc2c3ccc(OCCCCCC)cc3n(-c3ccnc(-c4ccc(F)c(C#N)c4F)c3)c2c1 1 2 A-D-A $581.707$ $9.914$
tBuCz2 CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc2[nH]c3ccc(-n4c5ccc(C(C)(C)C)cc5c5cc(C(C)(C)C)ccc54)cc3c2c1 3 0 Multi-D/A $722.033$ $14.705$
26DCzPPy c1cc(-c2cccc(-c3cccc(-n4c5ccccc5c5ccccc54)c3)n2)cc(-n2c3ccccc3c3ccccc32)c1 2 0 2D-0A $561.688$ $10.610$
mCBP c1cc(-c2cccc(-n3c4ccccc4c4ccccc43)c2)cc(-n2c3ccccc3c3ccccc32)c1 2 0 2D-0A $484.602$ $9.548$
benzoxazole Cc1sc2ccccc2c1C1=C(c2c(C)sc3ccccc23)C(F)(F)C(F)(F)C1(F)F 0 0 Unknown $468.487$ $8.563$
dbG2TAZ c1ccc2c(c1)ccc1[nH]c3ccc4ccccc4c3c12 1 0 1D-0A $267.331$ $5.628$
V–TPA C=Cc1ccc(COc2cccc3c2c2ccccc2n3-c2ccc(N(c3ccc(-n4c5ccccc5c5ccccc54)cc3)c3ccc(-n4c5ccccc5c5ccccc54)cc3)cc2)cc1 4 0 Multi-D/A $873.072$ $16.670$
CCDC O=C(c1ccccc1)c1ccc(-n2c3ccc(-n4c5ccccc5c5ccccc54)cc3c3cc(-n4c5ccccc5c5ccccc54)ccc32)cc1 3 0 Multi-D/A $677.807$ $12.209$
PXZPDO O=C(C=C(O)c1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 0 Multi-D/A $586.647$ $10.619$
DMAC-BP CC1(C)c2ccccc2N(c2ccc(C(=O)c3ccc(N4c5ccccc5C(C)(C)c5ccccc54)cc3)cc2)c2ccccc21 2 0 2D-0A $596.774$ $11.136$
SBTF c1ccc2c(c1)-c1ccccc1C21c2c(ccc3cc(-c4ccc5ccccc5c4)ccc23)-c2ccc(-c3cc4ccccc4c4ccccc34)c3cccc1c23 0 0 Unknown $718.899$ $15.130$
3-DPM N#CC(C#N)=C(c1cccc(N(c2ccccc2)c2ccccc2)c1)c1cccc(N(c2ccccc2)c2ccccc2)c1 2 2 2D-2A $564.692$ $10.475$
4-DPM N#CC(C#N)=C(c1ccc(N(c2ccccc2)c2ccccc2)cc1)c1ccc(N(c2ccccc2)c2ccccc2)cc1 2 2 2D-2A $564.692$ $10.475$
2CzPy c1ccc(-c2ccc(-n3c4ccccc4c4ccccc43)cc2-n2c3ccccc3c3ccccc32)nc1 2 0 2D-0A $485.590$ $8.943$
2PBCzPy c1ccc(-n2c3ccccc3c3cc(-c4ccc5c(c4)c4ccccc4n5-c4ccc(-c5ccccn5)c(-n5c6ccccc6c6cc(-c7ccc8c(c7)c7ccccc7n8-c7ccccc7)ccc65)c4)ccc32)cc1 4 0 Multi-D/A $968.176$ $18.471$
2BFCzPy c1ccc(-c2ccc(-n3c4ccccc4c4c5oc6ccccc6c5ccc43)cc2-n2c3ccccc3c3c4oc5ccccc5c4ccc32)nc1 2 0 2D-0A $665.752$ $12.742$
CNTPA-CZ N#Cc1ccc(N(c2ccccc2)c2cc(-n3c4ccccc4c4ccccc43)cc(-n3c4ccccc4c4ccccc43)c2)cc1 3 2 Multi-D/A $600.725$ $11.222$
CNTPA-PXZ N#Cc1ccc(N(c2ccccc2)c2cc(N3c4ccccc4Oc4ccccc43)cc(N3c4ccccc4Oc4ccccc43)c2)cc1 5 2 Multi-D/A $632.723$ $12.179$
BMZ-TZ c1cc(-c2nc3ccccc3[nH]2)cc(-c2nnnc(-c3cccc(-c4nc5ccccc5[nH]4)c3)c2-c2cccc(-c3nc4ccccc4[nH]3)c2)c1 0 0 Unknown $657.741$ $9.502$
TPAPm N#Cc1c(-c2ccc(N(c3ccccc3)c3ccccc3)cc2)ncnc1-c1ccc(N(c2ccccc2)c2ccccc2)cc1 2 2 2D-2A $591.718$ $10.622$
CzPm N#Cc1c(-c2ccc(-n3c4ccccc4c4ccccc43)cc2)ncnc1-c1ccc(-n2c3ccccc3c3ccccc32)cc1 2 2 2D-2A $587.686$ $9.876$
TCPZ C1=C(c2cccc(-n3c4ccccc4c4ccccc43)c2)NN(c2cccc(-n3c4ccccc4c4ccccc43)c2)N=C1c1cccc(-n2c3ccccc3c3ccccc32)c1 3 0 Multi-D/A $806.973$ $13.748$
TrisCz-Trz c1ccc(-n2c3ccccc3c3ccccc32)c(-c2nc(-c3ccccc3-n3c4ccccc4c4ccccc43)nc(-c3ccccc3-n3c4ccccc4c4ccccc43)n2)c1 3 1 Multi-D/A $804.957$ $14.164$
PTZ–TAZ c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2nnc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)o2)cc1 4 0 Multi-D/A $584.635$ $10.555$
DDMA-TXO2 CC1(C)c2ccccc2N(c2ccc3c(c2)C(C)(C)c2cc(N4c5ccccc5C(C)(C)c5ccccc54)ccc2S3(=O)=O)c2ccccc21 2 0 2D-0A $672.894$ $11.377$
UGH3 c1ccc([Si](c2ccccc2)(c2ccccc2)c2cccc([Si](c3ccccc3)(c3ccccc3)c3ccccc3)c2)cc1 0 0 Unknown $594.906$ $4.441$
V.2 Data availability
All computational data, analysis scripts, and molecular structures are available at https://github.com/TchapetNjafa.
Acknowledgments
We gratefully acknowledge Dr. Benjamin Panebei Samafou for his generous support and provision of computational resources, which were instrumental in enabling the numerical simulations and data analyses presented in this work.